### v2

# On the rates of convergence for learning with convolutional neural networks

Abstract
We study approximation and learning capacities of convolutional neural networks (CNNs) with one-side zero-padding and multiple channels. Our first result proves a new approximation bound for CNNs with certain constraint on the weights. Our second result gives new analysis on the covering number of feed-forward neural networks with CNNs as special cases. The analysis carefully takes into account the size of the weights and hence gives better bounds than the existing literature in some situations. Using these two results, we are able to derive rates of convergence for estimators based on CNNs in many learning problems. In particular, we establish minimax optimal convergence rates of the least squares based on CNNs for learning smooth functions in the nonparametric regression setting. For binary classification, we derive convergence rates for CNN classifiers with hinge loss and logistic loss. It is also shown that the obtained rates for classification are minimax optimal in some common settings.
Keywords: convolutional neural network, convergence rate, regression, classification, approximation, covering number
1 Introduction
Deep leaning has made remarkable successes in many applications and research fields such as image classification, speech recognition, natural language processing and scientific computing [LeCun et al., 2015; Goodfellow et al., 2016]. This breakthrough of deep learning also motivated many theoretical researches on understanding and explaining the empirical successes of deep neural networks from various perspectives. In particular, recent studies have established optimal approximations of smooth function classes by fully connected neural networks [Yarotsky, 2017, 2018; Shen et al., 2020; Lu et al., 2021]. It has also been showed that these networks can achieve minimax optimal rates of convergence in many learning problems, including nonparametric regression [Schmidt-Hieber, 2020; Kohler and Langer, 2021] and classification [Kim et al., 2021].
Our main interests in this paper are the approximation and learning properties of convolutional neural networks (CNNs), which are widely used in image classification and related applications [Krizhevsky et al., 2012]. Recently, substantial progress has been made in the theoretical study of CNNs. It has been shown that CNNs are universal for approximation [Zhou, 2020b] and universally consistent for nonparametric regression [Lin et al., 2022]. Approximation bounds and representational advantages of CNNs have been proven in several works [Oono and Suzuki, 2019; Zhou, 2020a; Fang et al., 2020; Mao et al., 2021]. Furthermore, rates of convergence of estimators based on CNNs were established for nonparametric regression [Zhou and Huo, 2024; Yang and Zhou, 2024] and classification [Kohler and Langer, 2020; Liu et al., 2021; Feng et al., 2023]. However, in contrast with the minimax optimal learning rates for fully connected neural networks, many of these results for CNNs are not optimal.
In this paper, we take a step to close this gap by providing new analysis on the approximation and learning capacities of CNNs. Specifically, we prove new bounds for the approximation of smooth functions by CNNs with certain constraint on the weights. We also derive bounds for the covering numbers of these networks. Using the obtained bounds, we are able to establish convergence rates for CNNs in many learning problems. These rates are known to be minimax optimal in several settings. We summarize our contributions in the following.
  1. (1)
    We prove the rates $\mathcal{O}(L^{-\alpha/d})$ for the approximation of smooth functions with smoothness $\alpha<(d+3)/2$ by CNNs, where $d$ is the dimension and $L$ is the depth of CNNs. The main advantage of our result is that we have an explicit control on the network weights (through $\kappa(\theta)$ defined by (2.3) below). It is proven that one can choose $\kappa(\theta)\leq M$ with $M=\mathcal{O}(L^{\frac{3d+3-2\alpha}{2d}})$ to ensure that the approximation rate $\mathcal{O}(L^{-\alpha/d})$ holds.
  2. (2)
    We provide a new framework to estimate the covering numbers of feed-forward neural networks. An application of this result gives the bound $\mathcal{O}(L\log(LM/\epsilon))$ for the $\epsilon$-covering number of CNNs with depth $L$ and weight constraint $\kappa(\theta)\leq M$. When $M$ grows at most polynomially on $L$, our bound is better than the general bound $\mathcal{O}(L^{2}\log(L/\epsilon))$ in the literature.
  3. (3)
    For regression, we establish the minimax optimal rate for the least squares regression with CNNs, when the regression function is smooth.
  4. (4)
    For binary classification, we establish rates of convergence for CNN classifiers with hinge loss and logistic loss, under the Tsybakov noise condition (4.3). For the hinge loss, the obtained rate for the excess classification risk is minimax optimal. For the logistic loss, the obtained rate may not be optimal for the excess classification risk. But it is optimal for the excess logistic risk, at least in some situations.
The remainder of this paper is organized as follows. In Section 2, we describe the architecture of convolutional neural networks used in this paper, and derive bounds for the approximation capacity and covering number of these networks. Sections 3 and 4 study the nonparametric regression and classification problems, respectively. We derive convergence rates of the excess risk for CNNs in these two sections. Section 5 concludes this paper with a discussion on future studies.
1.1 Notations
[…]
2 Convolutional neural networks
Let us first define convolutional neural networks used in this paper. Let ${\boldsymbol{w}}=(w_{1},\dots,w_{s})^{\intercal}\in\mathbb{R}^{s}$ be a filter with filter size $s\in[d]$. We define the convolution matrix $T_{\boldsymbol{w}}$ on $\mathbb{R}^{d}$ by
$T_{\boldsymbol{w}}:=\begin{pmatrix}w_{1}&\cdots&w_{s-1}&w_{s}\\ &\ddots&\ddots&\ddots&\ddots\\ &&w_{1}&\cdots&w_{s-1}&w_{s}\\ &&&w_{1}&\cdots&w_{s-1}\\ &&&&\ddots&\vdots\\ &&&&&w_{1}\end{pmatrix}\in\mathbb{R}^{d\times d}$
This convolution matrix corresponds to the one-sided padding and stride-one convolution by the filter ${\boldsymbol{w}}$. It is essentially the same as the convolution used in [Oono and Suzuki, 2019] (up to a matrix transpose). But it is different from the convolution matrix used in [Zhou, 2020a, b; Fang et al., 2020; Mao et al., 2021; Feng et al., 2023; Yang and Zhou, 2024], which is of dimension $(d+s)\times d$, rather than $d\times d$. So, in their setting, the network width increases after every application of the convolution with only one channel, while the network width remains the same with more channels in our setting. We define convolutional layers as follows. Let $s,J,J^{\prime}\in\mathbb{N}$ be a filter size, input channel size, and output channel size. For a filter ${\boldsymbol{w}}=(w_{i,j^{\prime},j})_{i\in[s],j^{\prime}\in[J^{\prime}],j\in[J]}\in\mathbb{R}^{s\times J^{\prime}\times J}$ and a bias vector ${\boldsymbol{b}}=(b_{1},\dots,b_{J^{\prime}})^{\intercal}\in\mathbb{R}^{J^{\prime}}$, we define the convolutional layer as an operator $\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}:\mathbb{R}^{d\times J}\to\mathbb{R}^{d\times J^{\prime}}$ by
$(\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}}))_{:,j^{\prime}}:=\sum_{j=1}^{J}T_{w_{:,j^{\prime},j}}x_{:,j}+b_{j^{\prime}},\quad{\boldsymbol{x}}=(x_{i,j})_{i\in[d],j\in[J]}\in\mathbb{R}^{d\times J},$
where we use “$+b_{j^{\prime}}$” to denote the vector addition “$+(b_{j^{\prime}},\dots,b_{j^{\prime}})^{\intercal}$” when there is no confusion. Next, we define convolutional neural networks. Let $s\in[d]$ and $J,L\in\mathbb{N}$ be the filter size, channel size and depth. We denote by $\mathcal{CNN}(s,J,L)$ the set of functions $f_{\boldsymbol{\theta}}$ that can be parameterized by ${\boldsymbol{\theta}}=({\boldsymbol{w}}^{(0)},{\boldsymbol{b}}^{(0)},\dots,{\boldsymbol{w}}^{(L-1)},{\boldsymbol{b}}^{(L-1)},{\boldsymbol{w}}^{(L)})$ in the following form
$f_{\boldsymbol{\theta}}({\boldsymbol{x}}):=\left\langle{\boldsymbol{w}}^{(L)},\sigma\circ\,{\rm Conv}\,_{{\boldsymbol{w}}^{(L-1)},{\boldsymbol{b}}^{(L-1)}}\circ\cdots\circ\sigma\circ\,{\rm Conv}\,_{{\boldsymbol{w}}^{(0)},{\boldsymbol{b}}^{(0)}}({\boldsymbol{x}})\right\rangle,\quad{\boldsymbol{x}}\in[0,1]^{d},$ (2.1)
where ${\boldsymbol{w}}^{(0)}\in\mathbb{R}^{s\times J\times 1},{\boldsymbol{b}}^{(0)}\in\mathbb{R}^{J},{\boldsymbol{w}}^{(L)}\in\mathbb{R}^{d\times J},{\boldsymbol{w}}^{(\ell)}\in\mathbb{R}^{s\times J\times J},{\boldsymbol{b}}^{(\ell)}\in\mathbb{R}^{J}$ for $\ell\in[L-1]$, and the activation $\sigma(t)=t\lor 0$ is the ReLU activation function. Note that we have assumed the channel sizes in each layers are the same, because we can always increase the channel sizes by adding appropriate zero filters and biases. For convenience, we will often view ${\boldsymbol{w}}^{(0)}\in\mathbb{R}^{s\times J\times J}$ by adding zeros to the filter and the input. The number of parameters in the network is $(sJ+1)JL+(d+s-sJ)J\lesssim J^{2}L$, which grows linearly on the depth $L$.
In order to control the complexity of convolutional neural networks, we introduce the following norm for the pair $({\boldsymbol{w}},{\boldsymbol{b}})\in\mathbb{R}^{s\times J\times J}\times\mathbb{R}^{J}$
$\|({\boldsymbol{w}},{\boldsymbol{b}})\|:=\max_{j^{\prime}\in[J]}\left(\left\|w_{:,j^{\prime},:}\right\|_{1}+|b_{j^{\prime}}|\right).$
Note that $\|({\boldsymbol{w}},{\boldsymbol{b}})\|$ quantifies the size of the affine transform $\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}$:
$\left\|\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}})\right\|_{\infty}$ $\leq\max_{j^{\prime}\in[J]}\left(\sum_{j=1}^{J}\left\|T_{w_{:,j^{\prime},j}}x_{:,j}\right\|_{\infty}+|b_{j^{\prime}}|\right)$
$\leq\|({\boldsymbol{w}},{\boldsymbol{b}})\|(\|{\boldsymbol{x}}\|_{\infty}\lor 1).$ (2.2)
Following the idea of [Jiao et al., 2023], we define a constraint on the weights as follows
$\kappa({\boldsymbol{\theta}}):=\|{\boldsymbol{w}}^{(L)}\|_{1}\prod_{\ell=0}^{L-1}\left(\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|\lor 1\right).$ (2.3)
For any $M\geq 0$, we denote the function class consisting of wights constrained CNNs by
$\mathcal{CNN}(s,J,L,M):=\left\{f_{\boldsymbol{\theta}}\in\mathcal{CNN}(s,J,L):\kappa({\boldsymbol{\theta}})\leq M\right\}.$
Several properties of this function class are summarized in Appendix A.1. In Sections 2.1 and 2.2, we study the approximation capacity and covering number of $\mathcal{CNN}(s,J,L,M)$. These results are used in Sections 3 and 4 to study the convergence rates of CNNs on the nonparametric regression and classification problems.
2.1 Approximation
We consider the capacity of CNNs for approximating smooth functions. Given a smoothness index $\alpha>0$, we write $\alpha=r+\beta$ where $r\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ and $\beta\in(0,1]$. Let $C^{r,\beta}(\mathbb{R}^{d})$ be the Hölder space with the norm
$\|f\|_{C^{r,\beta}(\mathbb{R}^{d})}:=\max\left\{\|f\|_{C^{r}(\mathbb{R}^{d})},\max_{\|{\boldsymbol{s}}\|_{1}=r}|\partial^{\boldsymbol{s}}f|_{C^{0,\beta}(\mathbb{R}^{d})}\right\},$
where ${\boldsymbol{s}}=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index and
$\|f\|_{C^{r}(\mathbb{R}^{d})}$ $:=\max_{\|{\boldsymbol{s}}\|_{1}\leq r}\|\partial^{\boldsymbol{s}}f\|_{L^{\infty}(\mathbb{R}^{d})},$
$|f|_{C^{0,\beta}(\mathbb{R}^{d})}$ $:=\sup_{{\boldsymbol{x}}\neq{\boldsymbol{y}}\in\mathbb{R}^{d}}\frac{|f({\boldsymbol{x}})-f({\boldsymbol{y}})|}{\|{\boldsymbol{x}}-{\boldsymbol{y}}\|_{2}^{\beta}}.$
Here, we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm since we only consider continuous functions. We write $C^{r,\beta}([0,1]^{d})$ for the Banach space of all restrictions to $[0,1]^{d}$ of functions in $C^{r,\beta}(\mathbb{R}^{d})$. The norm is given by $\|f\|_{C^{r,\beta}([0,1]^{d})}=\inf\{\|g\|_{C^{r,\beta}(\mathbb{R}^{d})}:g\in C^{r,\beta}(\mathbb{R}^{d})\mbox{ and }g=f\mbox{ on }[0,1]^{d}\}$. For convenience, we will denote the ball of $C^{r,\beta}([0,1]^{d})$ with radius $R>0$ by
$\mathcal{H}^{\alpha}(R):=\left\{f\in C^{r,\beta}([0,1]^{d}):\|f\|_{C^{r,\beta}([0,1]^{d})}\leq R\right\}.$
Note that, for $\alpha=1$, $\mathcal{H}^{1}(R)$ is a class of Lipschitz continuous functions.
Our first result estimates the error of approximating Hölder functions by CNNs.
Theorem 2.1.
Let $20<\alpha<(d+3)/$ 2 and $s\in[2:d]$. If $L\geq\lceil\frac{d-1}{s-1}\rceil$ and $M\gtrsim L^{\frac{3d+3-2\alpha}{2d}}$, then
$\sup_{h\in\mathcal{H}^{\alpha}(1)}\inf_{f\in\mathcal{CNN}(s,6,L,M)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim L^{-\frac{\alpha}{d}}.$
The approximation rate $\mathcal{O}(L^{-\alpha/d})$ is slightly better than $\mathcal{O}((L/\log L)^{-\alpha/d})$ in [Oono and Suzuki, 2019, Corollary 4] for ResNet-type CNNs. Furthermore, the result of [Oono and Suzuki, 2019] requires that the depth of residual blocks grows with the approximation error, while our result does not need any residual blocks. Our approximation rate is the same as the result of [Yang and Zhou, 2024], which used slightly different CNNs, and the rate in [Feng et al., 2023], which considered the approximation of smooth functions on spheres. The main advantage of our result is that Theorem 2.1 also provides an estimate on the weight constraint $M$, which is useful for the study of statistical properties of CNNs (see Sections 3 and 4).
The proof of Theorem 2.1 is given in Appendix A.3. Similar to many existing works for CNNs, such as [Oono and Suzuki, 2019; Zhou, 2020b], our proof uses the idea that smooth functions are well approximated by fully-connected neural networks and one can construct CNNs to implement fully-connected neural networks. Our result is based on the approximation bound for shallow neural networks proven in [Yang and Zhou, 2024]. To introduce the idea, let us denote the function class of shallow neural networks by
$\mathcal{NN}(N,M):=\left\{f({\boldsymbol{x}})=\sum_{i=1}^{N}c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i}):\sum_{i=1}^{N}|c_{i}|(\|{\boldsymbol{a}}_{i}\|_{1}+|b_{i}|)\leq M\right\}.$ (2.4)
It was shown by [Yang and Zhou, 2024, Corollary 2.4] that, if $\alpha<(d+3)/2$, then
$\sup_{h\in\mathcal{H}^{\alpha}(1)}\inf_{f\in\mathcal{NN}(N,M)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim N^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+3-2\alpha}}.$ (2.5)
In Appendix A.3, we show that functions in $\mathcal{NN}(N,M)$ can also be parameterized by a CNN with the same order of number of parameters. Specifically, for any $s\in[2:d]$ and $L_{0}=\lceil\frac{d-1}{s-1}\rceil$, we prove that
$\mathcal{NN}(N,M)\subseteq\mathcal{CNN}(s,6,NL_{0},3^{L_{0}+1}NM).$ (2.6)
Theorem 2.1 is a direct consequence of the approximation bound (2.5) and the inclusion (2.6).
The requirement on the smoothness $\alpha<(d+3)/2$ is of course due to the use of the approximation bound (2.5). For high smoothness $\alpha>(d+3)/2$, one can also derive approximation bounds for CNNs by using the results of [Yang and Zhou, 2024], as discussed in the following remark.
Remark 2.2.
If $\alpha>(d+3)/2$, it was shown by [Yang and Zhou, 2024, Theorem 2.1] that $\mathcal{H}^{\alpha}(1)\subseteq\mathcal{F}_{\sigma}(M)$ for some constant $M>0$, where
[…]
Here, $\mathbb{S}^{d}$ is the unit sphere of $\mathbb{R}^{d+1}$ and $\|\mu\|=|\mu|(\mathbb{S}^{d})$ is the total variation of the measure $\mu$. $\mathcal{F}_{\sigma}(M)$ is a ball with radius $M$ of the variation space corresponding to shallow ReLU neural networks studied in many recent papers, such as [Bach, 2017; E et al., 2022; Siegel and Xu, 2022, 2023; Siegel, 2023]. The function class $\mathcal{F}_{\sigma}(M)$ can be viewed as output functions of an infinitely wide neural network. It is the limit of $\mathcal{NN}(N,M)$ as the number of neurons $N\to\infty$ [Yang and Zhou, 2024, Proposition 2.2]. The recent work [Siegel, 2023] showed that
$\sup_{h\in\mathcal{F}_{\sigma}(1)}\inf_{f\in\mathcal{NN}(N,1)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim N^{-\frac{d+3}{2d}}.$
Combining this bound with the inclusion (2.6), we can obtain
$\sup_{h\in\mathcal{F}_{\sigma}(1)}\inf_{f\in\mathcal{CNN}(s,6,L,M)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim L^{-\frac{d+3}{2d}},$ (2.8)
for $M\gtrsim L$. This approximation bound can be used to study machine learning problems with smoothness assumption $\alpha>(d+3)/2$, see Remark 3.3 for example. But it seems that the bound (2.8) is not optimal for CNNs and high smoothness.
2.2 Covering number
In statistical learning theory, we often estimate generalization error of learning algorithms by certain complexities of models. The complexity we use in this paper is the covering number (or metric entropy) defined in the following.
Definition 2.3 (Covering number and entropy).
Let $\rho$ be a metric on a metric space $\mathcal{M}$ and $\mathcal{F}\subseteq\mathcal{M}$. For $\epsilon>0$, a set $\mathcal{S}\subseteq\mathcal{M}$ is called an $\epsilon$-cover (or $\epsilon$-net) of $\mathcal{F}$ if for any $x\in\mathcal{F}$, there exists $y\in\mathcal{S}$ such that $\rho(x,y)\leq\epsilon$. The $\epsilon$-covering number of $\mathcal{F}$ is defined by
$\mathcal{N}(\epsilon,\mathcal{F},\rho):=\min\{|\mathcal{S}|:\mathcal{S}\mbox{ is an $\epsilon$-cover of }\mathcal{F}\},$
where $|\mathcal{S}|$ is the cardinality of the set $\mathcal{S}$. The logarithm of the covering number $\log\mathcal{N}(\epsilon,\mathcal{F},\rho)$ is called (metric) entropy.
[…]
We consider neural networks of the following form
${\boldsymbol{f}}_{0}({\boldsymbol{x}})$ $={\boldsymbol{x}}\in[0,1]^{d},$ (2.9)
${\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})$ $=\sigma({\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))),\quad\ell\in[0:L-1],$
$f_{\boldsymbol{\theta}}({\boldsymbol{x}})$ $={\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{L}}({\boldsymbol{f}}_{L}({\boldsymbol{x}})),$
where ${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}:\mathbb{R}^{d_{\ell}}\to\mathbb{R}^{d_{\ell+1}}$ is an affine map parameterized by a vector ${\boldsymbol{\theta}}_{\ell}\in\mathbb{R}^{N_{\ell}}$ with $d_{0}=d$, $d_{L+1}=1$ and the vector of parameters ${\boldsymbol{\theta}}:=({\boldsymbol{\theta}}_{0}^{\intercal},\dots,{\boldsymbol{\theta}}_{L}^{\intercal})^{\intercal}\in\mathbb{R}^{N}$. Here, we use $N=\sum_{\ell=0}^{L}N_{\ell}$ to denote the number of parameters in the network. Note that we have restricted the input of the networks to $[0,1]^{d}$ for convenience. We assume that the parameterization satisfies the following conditions: for any ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in\mathbb{R}^{d_{\ell}}$ and $\ell\in[0:L]$,
$\|{\boldsymbol{\theta}}\|_{\infty}$ $\leq B,$ (2.10)
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})\|_{\infty}$ $\leq\gamma_{\ell}(\|{\boldsymbol{x}}\|_{\infty}\lor 1),$
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}}^{\prime})\|_{\infty}$ $\leq\gamma_{\ell}\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|_{\infty},$
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}}({\boldsymbol{x}})\|_{\infty}$ $\leq\lambda_{\ell}\|{\boldsymbol{\theta}}_{\ell}-{\boldsymbol{\theta}}_{\ell}^{\prime}\|_{\infty}(\|{\boldsymbol{x}}\|_{\infty}\lor 1),$
where ${\boldsymbol{\theta}}^{\prime}$ denotes any parameters satisfying $\|{\boldsymbol{\theta}}^{\prime}\|_{\infty}\leq B$.
Note that, if the affine map have the following matrix form
${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})={\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}}{\boldsymbol{x}}+{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}}=({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})\begin{pmatrix}{\boldsymbol{x}}\\ 1\end{pmatrix},$
then we can choose $\gamma_{\ell}$ to be the matrix operator norm (induced by $\|\cdot\|_{\infty}$)
$\gamma_{\ell}=\left\|({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})\right\|_{l^{\infty}\to l^{\infty}},$
and choose $\lambda_{\ell}$ to be the Lipschitz constant of the parameterization
$\left\|({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})-({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}})\right\|_{l^{\infty}\to l^{\infty}}\leq\lambda_{\ell}\|{\boldsymbol{\theta}}_{\ell}-{\boldsymbol{\theta}}_{\ell}^{\prime}\|_{\infty}.$
Recall that the matrix operator norm $\left\|({\boldsymbol{A}},{\boldsymbol{b}})\right\|_{l^{\infty}\to l^{\infty}}$ is the maximal $1$-norm of rows of the matrix $({\boldsymbol{A}},{\boldsymbol{b}})$. In our constructions, $({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})$ is linear on the parameter ${\boldsymbol{\theta}}_{\ell}$, which implies that we can choose $\lambda_{\ell}\lesssim d_{\ell}+1$.
The next lemma estimates the covering numbers of the neural networks described above.
Lemma 2.4.
Let $\mathcal{F}$ be the class of functions $f_{\boldsymbol{\theta}}$ that can be parameterized in the form (2.9), where the parameterization satisfies (2.10) with $\lambda_{\ell}\geq 0$ and $\gamma_{\ell}\geq 1$ for $\ell=[0:L]$. Then, the $\epsilon$-covering number of $\mathcal{F}$ in the $L^{\infty}([0,1]^{d})$ norm satisfies
$\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{L^{\infty}([0,1]^{d})})\leq(C_{L}B/\epsilon)^{N},$
where $N$ is the number of parameters and $C_{L}$ can be computed inductively by
$C_{0}=\lambda_{0},\quad C_{\ell+1}=\gamma_{\ell+1}C_{\ell}+\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}.$
In particular,
$C_{L}\leq\left(\sum_{j=0}^{L}\lambda_{j}\right)\prod_{i=0}^{L}\gamma_{i}.$
Proof.
For any ${\boldsymbol{\theta}},{\boldsymbol{\theta}}^{\prime}\in[-B,B]^{N}$ with $\|{\boldsymbol{\theta}}-{\boldsymbol{\theta}}^{\prime}\|_{\infty}\leq\epsilon$, we claim that, for any ${\boldsymbol{x}}\in[0,1]^{d}$ and $\ell\in[0:L]$,
$\|{\boldsymbol{f}}_{\ell}({\boldsymbol{x}})\|_{\infty}\leq\prod_{i=-1}^{\ell-1}\gamma_{i},$
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}}({\boldsymbol{f}}_{\ell}^{\prime}({\boldsymbol{x}}))\|_{\infty}\leq C_{\ell}\epsilon,$
$C_{\ell}\leq\left(\sum_{j=0}^{\ell}\lambda_{j}\right)\prod_{i=0}^{\ell}\gamma_{i},$
where we set $\gamma_{-1}=1$ and ${\boldsymbol{f}}_{\ell}^{\prime}$ denotes the function in (2.9) parameterized by ${\boldsymbol{\theta}}^{\prime}$. Thus, any $\epsilon$-cover of $[-B,B]^{N}$ gives a $C_{L}\epsilon$-cover of $\mathcal{F}$ in the $L^{\infty}([0,1]^{d})$ norm. Since the $\epsilon$-covering number of $[-B,B]^{N}$ is at most $(B/\epsilon)^{N}$, we get the desire bound for $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{L^{\infty}([0,1]^{d})})$.
We prove the claim by induction on $\ell\in[0:L]$. The claim is trivial for $\ell=0$ by definition. Assume that the claim is true for some $0\leq\ell<L$, we are going to prove it for $\ell+1$. By induction hypothesis,
$\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})\|_{\infty}\leq\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))\|_{\infty}\leq\gamma_{\ell}(\|{\boldsymbol{f}}_{\ell}({\boldsymbol{x}})\|_{\infty}\lor 1)\leq\prod_{i=0}^{\ell}\gamma_{i},$
where we used $\gamma_{i}\geq 1$ in the last inequality. By the Lipschitz continuity of ReLU,
$\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})-{\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}})\|_{\infty}\leq\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}}({\boldsymbol{f}}_{\ell}^{\prime}({\boldsymbol{x}}))\|_{\infty}\leq C_{\ell}\epsilon.$
Therefore,
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}}))\|_{\infty}$
$\leq$ $\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))\|_{\infty}+\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}}))\|_{\infty}$
$\leq$ $\lambda_{\ell+1}\epsilon(\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})\|_{\infty}\lor 1)+\gamma_{\ell+1}\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})-{\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}})\|_{\infty}$
$\leq$ $\left(\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}+\gamma_{\ell+1}C_{\ell}\right)\epsilon=C_{\ell+1}\epsilon.$
Finally, by induction hypothesis and $\gamma_{\ell+1}\geq 1$,
$C_{\ell+1}$ $=\gamma_{\ell+1}C_{\ell}+\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}$
$\leq\left(\sum_{j=0}^{\ell}\lambda_{j}\right)\prod_{i=0}^{\ell+1}\gamma_{i}+\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}$
$\leq\left(\sum_{j=0}^{\ell+1}\lambda_{j}\right)\prod_{i=0}^{\ell+1}\gamma_{i},$
which completes the proof. ∎
Now, we apply Lemma 2.4 to the convolutional neural network $\mathcal{CNN}(s,J,L,M)$. In this case, ${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}=\,{\rm Conv}\,_{{\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)}}$ for $\ell\in[0:L-1]$ and ${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{L}}(\cdot)=\langle{\boldsymbol{w}}^{(L)},\cdot\rangle$. By Proposition A.2 in the Appendix, we can assume $\|{\boldsymbol{w}}^{(L)}\|_{1}\leq M$ and $\|{\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|\leq 1$ for all $\ell\in[0:L-1]$, which implies $B=M\lor 1$. Using the inequality (2.2) and
$\left\|\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}})-\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}}^{\prime})\right\|_{\infty}$
$\leq$ $\max_{j^{\prime}\in[J]}\left(\sum_{j=1}^{J}\left\|T_{w_{:,j^{\prime},j}}x_{:,j}-T_{w_{:,j^{\prime},j}}x_{:,j}^{\prime}\right\|_{\infty}\right)$
$\leq$ $\|({\boldsymbol{w}},{\boldsymbol{b}})\|\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|_{\infty},$
we can set $\gamma_{\ell}=1$ and $\gamma_{L}=M$. It is easy to see that we can choose $\lambda_{\ell}=sJ+1$ and $\lambda_{L}=dJ$. Consequently,
$C_{L}\leq\left(\sum_{j=0}^{L}\lambda_{j}\right)\prod_{i=0}^{L}\gamma_{i}=(dJ+sJL+L)M\leq 3dJLM,$
where we use $s\leq d$ in the last inequality. We summarize the result in the next theorem.
Theorem 2.5.
Let $s,J,L\in\mathbb{N}$ and $M\geq 1$. The entropy of $\mathcal{CNN}(s,J,L,M)$ satisfies
$\log\mathcal{N}(\epsilon,\mathcal{CNN}(s,J,L,M),\|\cdot\|_{L^{\infty}([0,1]^{d})})\leq N\log(3dJLM^{2}/\epsilon),$
where $N=(sJ+1)JL+(d+s-sJ)J$ is the number of parameters in the network.
In the analysis of neural networks, many papers, such as [Schmidt-Hieber, 2020; Feng et al., 2023], simply assume that the parameters in the networks are bounded. In this case, the entropy is often bounded as $\mathcal{O}(NL\log(N/\epsilon))$, where $N$ is the number of parameters and $L$ is the depth. For convolutional neural networks with bounded width, we have $N\asymp L$ and hence the entropy would be $\mathcal{O}(L^{2}\log(L/\epsilon))$. For comparison, Theorem 2.5 gives a bound $\mathcal{O}(L\log(LM/\epsilon))$. If one only assumes that the parameters are bounded by $B$, then $M\lesssim B^{L}$ and our bound is consistent with the previous bound. However, if the weight constraint $M$ grows at most polynomially on $L$, then we get a better bound $\mathcal{O}(L\log(L/\epsilon))$ on the entropy. This improvement is essential to obtain optimal rates for many learning algorithms that we discuss in next two sections.
3 Regression
In this section, we consider the classical nonparametric regression problem. Assume that $({\boldsymbol{X}},Y)$ is a $[0,1]^{d}\times\mathbb{R}$-valued random vector satisfying $\mathbb{E}[Y^{2}]<\infty$. Let us denote the marginal distribution of ${\boldsymbol{X}}$ by $\mu$ and the regression function by
$h({\boldsymbol{x}}):=\mathbb{E}[Y|{\boldsymbol{X}}={\boldsymbol{x}}].$
Suppose we are given a data set of $n$ samples $\mathcal{D}_{n}=\{({\boldsymbol{X}}_{i},Y_{i})\}_{i=1}^{n}$, which are independent and have the same distribution as the random vector $({\boldsymbol{X}},Y)$. The goal of nonparametric regression problem is to construct an estimator $\widehat{f}_{n}$, based on $\mathcal{D}_{n}$, to reconstruct the regression function $h$. The estimation performance is evaluated by the $L^{2}$-error
$\|\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}=\mathbb{E}_{\boldsymbol{X}}\left[(\widehat{f}_{n}({\boldsymbol{X}})-h({\boldsymbol{X}}))^{2}\right].$
One of the popular algorithms to solve the regression problem is the empirical least squares
$\widehat{f}_{n}\in\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}(f({\boldsymbol{X}}_{i})-Y_{i})^{2},$ (3.1)
where $\mathcal{F}_{n}$ is a prescribed hypothesis class. For simplicity, we assume here and in the sequel that the minimum above indeed exists. We are interested in the case that the function class $\mathcal{F}_{n}$ is parameterized by a CNN. In order to study the convergence rate of $\widehat{f}_{n}\to h$ as $n\to\infty$, we will assume that $h\in\mathcal{H}^{\alpha}(R)$ for some constant $R>0$ and make the following assumption on the distribution of $({\boldsymbol{X}},Y)$: there exists a constant $c>0$ such that
$\mathbb{E}\left[\exp(cY^{2})\right]<\infty.$ (3.2)
In statistical analysis of learning algorithms, we often require that the hypothesis class is uniformly bounded. We define the truncation operator $\pi_{B}$ with level $B>0$ for real-valued functions $f$ as
$\pi_{B}f({\boldsymbol{x}})=\begin{cases}B\quad&f({\boldsymbol{x}})>B,\\ f({\boldsymbol{x}})\quad&|f({\boldsymbol{x}})|\leq B,\\ -B\quad&f({\boldsymbol{x}})<-B.\end{cases}$ (3.3)
Note that the truncation operator can be implemented by a CNN (see Lemma A.7 for example). Since we assume that the regression function $h$ is bounded, truncating the output of the estimator $\widehat{f}_{n}$ appropriately dose not increase the estimation error. The following theorem provides convergence rates for least squares estimators based on CNNs.
Theorem 3.1.
Assume that the condition (3.2) holds and the regression function $h\in\mathcal{H}^{\alpha}(R)$ for some $20<\alpha<(d+3)/$ 2 and $R>0$. Let $\widehat{f}_{n}$ be the estimator defined by (3.1) with $\mathcal{F}_{n}=\mathcal{CNN}(s,J,L_{n},M_{n})$, where $s\in[2:d]$, $J\geq 6$ and
$L_{n}\asymp\left(\frac{n}{\log^{3}n}\right)^{\frac{d}{2\alpha+d}},\quad\left(\frac{n}{\log^{3}n}\right)^{\frac{3d+3-2\alpha}{4\alpha+2d}}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n).$
If $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$, then
$\mathbb{E}_{\mathcal{D}_{n}}\left[\|\pi_{B_{n}}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]\lesssim\left(\frac{\log^{3}n}{n}\right)^{\frac{2\alpha}{2\alpha+d}}.$
It is well-known that the rate $n^{-\frac{2\alpha}{2\alpha+d}}$ is minimax optimal for learning functions in $\mathcal{H}^{\alpha}(R)$ [Stone, 1982]:
$\inf_{\widehat{f}_{n}}\sup_{h\in\mathcal{H}^{\alpha}(R)}\mathbb{E}_{\mathcal{D}_{n}}\left[\|\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]\gtrsim n^{-\frac{2\alpha}{2\alpha+d}},$
where the infimum is taken over all estimators based on the training data $\mathcal{D}_{n}$. Recent works have established the minimax rates (up to logarithm factors) for least squares estimators using fully-connected neural networks [Schmidt-Hieber, 2020; Kohler and Langer, 2021; Yang and Zhou, 2023]. For convolutional neural networks, [Oono and Suzuki, 2019] proved the optimal rates for ResNet-type CNNs, under the requirement that the depth of the residual blocks grows with the sample size $n$, or the residual blocks are suitably masked. Theorem 3.1 removes the requirements on the residual blocks for low smoothness $\alpha<(d+3)/2$.
Our proof of Theorem 3.1 is based on the following lemma from [Kohler and Langer, 2021, Lemma 18, Appendix B]. It decomposes the estimation error of the estimator into generalization error and approximation error, and bounds the generalization error by the covering number of the hypothesis class $\mathcal{F}_{n}$.
Lemma 3.2.
Assume that the condition (3.2) holds. Let $\widehat{f}_{n}$ be the estimator (3.1) and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$. Then,
$\mathbb{E}_{\mathcal{D}_{n}}\left[\|\pi_{B_{n}}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]$
$\leq$ $\frac{c_{2}(\log n)^{2}\sup_{{\boldsymbol{X}}_{1:n}}\log(\mathcal{N}(n^{-1}B_{n}^{-1},\pi_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}({\boldsymbol{X}}_{1:n})})+1)}{n}+2\inf_{f\in\mathcal{F}_{n}}\|f-h\|_{L^{2}(\mu)}^{2},$
for $n>1$ and some constant $c_{2}>0$ (independent of $n$ and $\widehat{f}_{n}$), where ${\boldsymbol{X}}_{1:n}=({\boldsymbol{X}}_{1},\dots,{\boldsymbol{X}}_{n})$ denotes a sequence of sample points in $[0,1]^{d}$ and $\mathcal{N}(\epsilon,\pi_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$ is the $\epsilon$-covering number of the function class $\pi_{B_{n}}\mathcal{F}_{n}:=\{\pi_{B_{n}}f,f\in\mathcal{F}_{n}\}$ in the metric $\|f-g\|_{L^{1}({\boldsymbol{X}}_{1:n})}=\frac{1}{n}\sum_{i=1}^{n}|f({\boldsymbol{X}}_{i})-g({\boldsymbol{X}}_{i})|$.
We now prove Theorem 3.1 by using Theorem 2.1 to bound the approximation error and using Theorem 2.5 to estimate the covering number.
Proof of Theorem 3.1.
It is easy to see that $\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\leq\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}([0,1]^{d})})$. Notice that the projection $\pi_{B}$ does not increase the covering number. By Theorem 2.5, we have
$\log\mathcal{N}(\epsilon,\pi_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{\infty}([0,1]^{d})})\lesssim L_{n}\log(L_{n}M_{n}/\epsilon).$
If $M_{n}\gtrsim L_{n}^{\frac{3d+3-2\alpha}{2d}}$, by Theorem 2.1, we get
$\inf_{f\in\mathcal{F}_{n}}\|f-h\|_{L^{2}(\mu)}^{2}\lesssim L_{n}^{-\frac{2\alpha}{d}}.$
As a consequence, Lemma 3.2 implies
$\mathbb{E}_{\mathcal{D}_{n}}\left[\|\pi_{B_{n}}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]$ $\lesssim\frac{\log^{2}n}{n}L_{n}\log(nL_{n}M_{n}B_{n})+L_{n}^{-\frac{2\alpha}{d}}$
$\lesssim\frac{\log^{3}n}{n}L_{n}+L_{n}^{-\frac{2\alpha}{d}},$
where we use $L_{n},M_{n}\lesssim\,{\rm Poly}\,(n)$ and $B_{n}=c\log n$ in the last inequality. Finally, by choosing $L_{n}\asymp(n/\log^{3}n)^{d/(2\alpha+d)}$, we finish the proof. ∎
Remark 3.3.
As we noted in Remark 2.2, if $\alpha>(d+3)/2$, then $\mathcal{H}^{\alpha}(1)\subseteq\mathcal{F}_{\sigma}(R)$ for some constant $R>0$. When the regression function $h\in\mathcal{F}_{\sigma}(R)$, we can use the approximation bound (2.8) to show that
$\mathbb{E}_{\mathcal{D}_{n}}\left[\|\pi_{B_{n}}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]\lesssim\left(\frac{\log^{3}n}{n}\right)^{\frac{d+3}{2d+3}},$
if we choose $L_{n}\asymp(n/\log^{3}n)^{d/(2d+3)}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n)$. This rate is minimax optimal (up to logarithm factors) for the function class $\mathcal{F}_{\sigma}(R)$ [Yang and Zhou, 2024]. For comparison, [Yang and Zhou, 2024] only established the sub-optimal rate $\mathcal{O}(n^{-\frac{d+3}{3d+3}}\log^{4}n)$ for CNNs. Our result is also better than the recent analysis of CNNs in [Zhou and Huo, 2024], which proved the rate $\mathcal{O}(n^{-1/3}\log^{2}n)$ for $\mathcal{H}^{\alpha}(R)$ with $\alpha>(d+4)/2$.
4 Binary classification
In binary classification, we observe a dataset $\mathcal{D}_{n}:=\{({\boldsymbol{X}}_{i},Y_{i}):i=1,\dots,n\}$ of $n$ i.i.d. copies of a random vector $({\boldsymbol{X}},Y)$, where we assume that the input vector ${\boldsymbol{X}}\in[0,1]^{d}$ and the label $Y\in\{-1,1\}$. The marginal distribution of ${\boldsymbol{X}}$ is denoted by $\mathbb{P}_{\boldsymbol{X}}$ and the conditional class probability function is denoted by
$\eta({\boldsymbol{x}}):=\mathbb{P}(Y=1|{\boldsymbol{X}}={\boldsymbol{x}}).$
For a real-valued function $f$ defined on $[0,1]^{d}$, we can define a classifier $\mathcal{C}_{f}({\boldsymbol{x}}):=\,{\rm sgn}\,(f({\boldsymbol{x}}))$. The classification error of $f$ is defined as
$\mathcal{E}(f)=\mathbb{E}_{{\boldsymbol{X}},Y}[\mathcal{C}_{f}({\boldsymbol{X}})\neq Y]=\mathbb{E}_{{\boldsymbol{X}},Y}[\boldsymbol{1}(Yf({\boldsymbol{X}})<0)],$
where $\boldsymbol{1}(\cdot)$ is $1$ if ()($\cdot$)( ⋅ ) is true, and is 00 otherwise. A Bayes classifier $\mathcal{C}^{*}=\mathcal{C}_{f^{*}}$ is a classifier that minimizes the classification error $\mathcal{E}(f^{*})=\min_{f\in\mathcal{M}}\mathcal{E}(f)$, where $\mathcal{M}$ is the set of all measurable functions on $[0,1]^{d}$. Note that $\mathcal{C}^{*}=\,{\rm sgn}\,(2\eta-1)$ is a Bayes classifier and $\mathcal{E}(\mathcal{C}^{*})=\frac{1}{2}\mathbb{E}[1-|2\eta-1|]$. The goal of binary classification is to construct a classifier with small classification error by using the dataset $\mathcal{D}_{n}$.
Since we only have finite observed samples, one natural approach to estimate the Bayes classifier is the empirical risk minimization (with $0-1$ loss)
$\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{1}(Y_{i}f({\boldsymbol{X}}_{i})<0),$ (4.1)
where $\mathcal{F}_{n}$ is a prescribed function class. However, this procedure is often computational infeasible due to the NP-hardness of the minimization problem. In general, one replaces the $0-1$ loss by surrogate losses. For a given surrogate loss function $\phi:\mathbb{R}\to[0,\infty)$, the $\phi$-risk is defined as
$\mathcal{L}_{\phi}(f)$ $:=\mathbb{E}_{{\boldsymbol{X}},Y}[\phi(Yf({\boldsymbol{X}}))]$
$=\mathbb{E}_{\boldsymbol{X}}[\eta({\boldsymbol{X}})\phi(f({\boldsymbol{X}}))+(1-\eta({\boldsymbol{X}}))\phi(-f({\boldsymbol{X}}))].$
Its minimizer is denoted by $f^{*}_{\phi}\in\operatorname*{argmin}_{f\in\mathcal{M}}\mathcal{L}_{\phi}(f)$. Note that $f^{*}_{\phi}$ can be explicitly computed by using the conditional class probability function $\eta$ for many convex loss functions $\phi$ [Zhang, 2004; Wu et al., 2007]. Instead of using (4.1), we can estimate the Bayes classifier by minimizing the empirical $\phi$-risk over a function class $\mathcal{F}_{n}$:
$\widehat{f}_{\phi,n}\in\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}\phi(Y_{i}f({\boldsymbol{X}}_{i})).$ (4.2)
The goal of this section is to estimate the convergence rates of the excess classification risk and excess $\phi$-risk defined by
$\mathcal{R}(\widehat{f}_{\phi,n})$ $:=\mathcal{E}(\widehat{f}_{\phi,n})-\mathcal{E}(\mathcal{C}^{*}),$
$\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})$ $:=\mathcal{L}_{\phi}(\widehat{f}_{\phi,n})-\mathcal{L}_{\phi}(f^{*}_{\phi}),$
when $\mathcal{F}_{n}$ is parameterized by a CNN. The convergence rates certainly depend on properties of the conditional class probability function $\eta$. One of the well known assumptions on $\eta$ is the Tsybakov noise condition [Mammen and Tsybakov, 1999; Tsybakov, 2004]: there exist $q\in[0,\infty]$ and $c_{q}>0$ such that for any $t>0$,
$\mathbb{P}_{\boldsymbol{X}}(|2\eta({\boldsymbol{X}})-1|\leq t)\leq c_{q}t^{q}.$ (4.3)
The constant $q$ is usually called the noise exponent. It is obvious that the Tsybakov noise condition always holds for $q=0$, whereas noise exponent $q=\infty$ means that $\eta$ is bounded away from the critical level $1/2$. We will consider classifications with hinge loss and logistic loss under the Tsybakov noise condition.
4.1 Hinge loss
For the hinge loss $\phi(t)=\max\{1-t,0\}$, we have $f^{*}_{\phi}=\,{\rm sgn}\,(2\eta-1)=\mathcal{C}^{*}$ and $\mathcal{L}_{\phi}(f^{*}_{\phi})=\mathbb{E}[1-|2\eta-1|]$. It is well known that the following calibration inequality holds [Zhang, 2004; Bartlett et al., 2006]
$\mathcal{R}(f)\leq\mathcal{R}_{\phi}(f).$ (4.4)
Hence, any convergence rate for the excess $\phi$-risk $\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})$ implies the same convergence rate for the excess classification risk $\mathcal{R}(\widehat{f}_{\phi,n})$. One can also check that [Zhang, 2004, Section 3.3], if $|f|\leq 1$, then
$\mathcal{R}_{\phi}(f)=\mathbb{E}[|f-f^{*}_{\phi}||2\eta-1|].$ (4.5)
To use this equality, it is natural to truncate the output of the estimator by using the truncation operator $\pi_{1}$ defined by (3.3).
In the following theorem, we provide convergence rates for the excess $\phi$-risk of the CNN classifier with hinge loss, under the assumption that the conditional class probability function $\eta$ is smooth and satisfies the Tsybakov noise condition.
Theorem 4.1.
Assume the noise condition (4.3) holds for some $q\in[0,\infty]$ and $\eta\in\mathcal{H}^{\alpha}(R)$ for some $20<\alpha<(d+3)/$ 2 and $R>0$. Let $\phi$ be the hinge loss and $\widehat{f}_{\phi,n}$ be the estimator defined by (4.2) with $\mathcal{F}_{n}=\{\pi_{1}f:f\in\mathcal{CNN}(s,J,L_{n},M_{n})\}$, where $s\in[2:d]$, $J\geq 6$ and
$L_{n}\asymp\left(\frac{n}{\log^{2}n}\right)^{\frac{d}{(q+2)\alpha+d}},\quad\left(\frac{n}{\log^{2}n}\right)^{\frac{3d+3}{2(q+2)\alpha+2d}}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n),$
then, for sufficiently large $n$,
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\right]\lesssim\left(\frac{\log^{2}n}{n}\right)^{\frac{(q+1)\alpha}{(q+2)\alpha+d}}.$
Audibert and Tsybakov [2007] showed that the minimax lower bound for the excess classification risk is
$\inf_{\widehat{f}_{n}}\sup_{\eta}\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}(\widehat{f}_{n})\right]\gtrsim n^{-\frac{(q+1)\alpha}{(q+2)\alpha+d}},$ (4.6)
where the supremum is taken over all $\eta\in\mathcal{H}^{\alpha}(R)$ that satisfies Tsybakov noise condition (4.3) and the infimum is taken over all estimators based on the training data $\mathcal{D}_{n}$. Hence, by the calibration inequality (4.4), the convergence rate in Theorem 4.1 is minimax optimal up to a logarithmic factor. Similar results have been established in [Kim et al., 2021] for fully connected neural networks with hinge loss. However, their results rely on the sparsity of neural networks and hence one need to optimize over different network architectures to obtain the optimal rate, which is hard to implement due to the unknown locations of the non-zero parameters. Our result show that CNNs, whose architecture is specifically defined, are able to achieve the optimal rate.
4.2 Logistic loss
For the logistic loss $\phi(t)=\log(1+e^{-t})$, we have $f^{*}_{\phi}=\log(\frac{\eta}{1-\eta})$ and $\mathcal{L}_{\phi}(f^{*}_{\phi})=\mathbb{E}[-\eta\log\eta-(1-\eta)\log(1-\eta)]$. Consequently, one can show that
$\mathcal{R}_{\phi}(f)=\mathbb{E}\left[\eta\log\left(\eta(1+e^{-f})\right)+(1-\eta)\log\left((1-\eta)(1+e^{f})\right)\right].$
Let us denote the KL-divergence by
$\mathcal{D}_{KL}(p,q):=p\log\left(\frac{p}{q}\right)+(1-p)\log\left(\frac{1-p}{1-q}\right),\quad p,q\in[0,1],$
where $\mathcal{D}_{KL}(p,q)=\infty$ if $q=0$ and $p\neq 0$, or $q=1$ and $p\neq 1$. If we define the logistic function by
$\psi(t):=\frac{1}{1+e^{-t}}\in[0,1],\quad t\in[-\infty,\infty],$ (4.7)
then a direct calculation shows that $\eta=\psi(f_{\phi}^{*})$ and
$\mathcal{R}_{\phi}(f)=\mathbb{E}[\mathcal{D}_{KL}(\eta,\psi(f))].$ (4.8)
When the Tsybakov noise condition (4.3) holds, we have the following calibration inequality [Steinwart and Christmann, 2008, Theorem 8.29]
$\mathcal{R}(f)\leq 4c_{q}^{\frac{1}{q+2}}\mathcal{R}_{\phi}(f)^{\frac{q+1}{q+2}}.$ (4.9)
For the logistic loss, the convergence rate depends not only on the Tsybakov noise condition, but also upon the Small Value Bound (SVB) condition introduced by [Bos and Schmidt-Hieber, 2022]. We say the distribution of $({\boldsymbol{X}},Y)$ satisfies the SVB condition, if there exists $\beta\geq 0$ and $C_{\beta}>0$ such that for any $t\in(0,1]$,
$\mathbb{P}_{\boldsymbol{X}}(\eta({\boldsymbol{X}})\leq t)\leq C_{\beta}t^{\beta},\quad\mathbb{P}_{\boldsymbol{X}}(1-\eta({\boldsymbol{X}})\leq t)\leq C_{\beta}t^{\beta}.$ (4.10)
Note that this condition always holds for $\beta=0$ with $C_{\beta}=1$. The index $\beta$ is completely determined by the behavior of $\eta$ near 00 and $1$. If $\eta$ is bounded away form 00 and $1$, then the SVB condition holds for all $\beta>0$. In contrast, the Tsybakov noise condition provides a control on the behavior of $\eta$ near the decision boundary $\{{\boldsymbol{x}}:\eta({\boldsymbol{x}})=1/2\}$. This difference is due to the loss: the $0-1$ loss only cares about the classification error, while the logistic loss measures how well the conditional class probability is estimated in the KL-divergence (4.8), which puts additional emphasis on small and large conditional class probabilities.
The following theorem gives convergence rates for CNNs under the SVB condition. As pointed out by [Bos and Schmidt-Hieber, 2022], we do not get any gain in the convergence rate when the SVB index $\beta>1$. So, we assume $\beta\in[0,1]$ in the theorem.
Theorem 4.2.
Assume the SVB condition (4.10) holds for some $\beta\in[0,1]$ and $\eta\in\mathcal{H}^{\alpha}(R)$ for some $20<\alpha<(d+3)/$ 2 and $R>0$. Let $\phi$ be the logistic loss and $\widehat{f}_{\phi,n}$ be the estimator defined by (4.2) with $\mathcal{F}_{n}=\{\pi_{B_{n}}f:f\in\mathcal{CNN}(s,J,L_{n},M_{n})\}$, where $s\in[2:d]$, $J\geq 6$ and
$L_{n}\asymp\left(\frac{n}{\log n}\right)^{\frac{d}{(1+\beta)\alpha+d}},\quad\left(\frac{n}{\log n}\right)^{\frac{3d+3+2\alpha}{2(1+\beta)\alpha+2d}}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n),\quad B_{n}\asymp\log n,$
then, for sufficiently large $n$,
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\right]\lesssim\left(\frac{\log n}{n}\right)^{\frac{(1+\beta)\alpha}{(1+\beta)\alpha+d}}\log n.$
The convergence rate in Theorem 4.2 is the same as [Bos and Schmidt-Hieber, 2022, Theorem 3.3], which studied multi-class classification using fully-connected deep neural networks with cross entropy loss. If, in addition, the Tsybakov noise condition (4.3) holds, by combining Theorem 4.2 with the calibration inequality (4.9), we can get the following convergence rate for the classification risk:
$\mathbb{E}_{\mathcal{D}_{n}}[\mathcal{R}(\widehat{f}_{\phi,n})]\lesssim n^{-\frac{q+1}{q+2}\frac{(1+\beta)\alpha}{(1+\beta)\alpha+d}}\log^{2}n.$
Note that this rate is the same as the optimal rate (4.6) up to a logarithmic factor, when $q=0$ and $\beta=1$. For $\beta=0$, the obtained rate is not minimax optimal for the excess classification risk. However, as shown by [Zhang et al., 2024, Corollary 2.1], the rate in Theorem 4.2 is indeed minimax optimal up to a logarithmic factor for the excess $\phi$-risk when $\beta=0$. So, even if the logistic classification can achieve the minimax optimal convergence rates for classification, it is in general not possible to derive it through the rates for excess $\phi$-risk.
There are other papers [Kohler and Langer, 2020; Liu et al., 2021] studying the convergence rates of CNNs with logistic loss. Kohler and Langer [2020] imposed a max-pooling structure for the conditional class probability that is related to the structure of convolutional networks. So, their result is not comparable to ours. Liu et al. [2021] used a similar setting as ours and derived the rate $n^{-\frac{\alpha}{2\alpha+2(\alpha\lor d)}}$ (ignoring logarithmic factors) for the excess $\phi$-risk under the assumption that $\eta$ is supported on a manifold of $d$ dimension. Our rate $n^{-\frac{\alpha}{\alpha+d}}$ in Theorem 4.2 (for $\beta=0$) is better, but their result can be applied to low-dimensional distributions. It would be interesting to generalize our result to these distributions.
4.3 Sketch of proofs
We sketch the proofs of Theorems 4.1 and 4.2 in this subsection and give the detailed proofs in the Appendix. Our proofs of Theorems 4.1 and 4.2 are based on the following lemma, which is summarized from [Kim et al., 2021, Appendix A.2].
Lemma 4.3.
Let $\phi$ be a surrogate loss function and $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ be a sequence of function classes. Assume that the random vector $({\boldsymbol{X}},Y)\in[0,1]^{d}\times\{-1,1\}$ and the following regularity conditions hold:
  1. (A1)
    $\phi$ is Lipschitz, i.e., there exists a constant $c_{1}>0$ such that $|\phi(t_{1})-\phi(t_{2})|\leq c_{1}|t_{1}-t_{2}|$ for any $t_{1},t_{2}\in\mathbb{R}$.
  2. (A2)
    There exist a positive sequence $\{a_{n}\}_{n\in\mathbb{N}}$ and $f_{n}\in\mathcal{F}_{n}$ such that
    $\mathcal{R}_{\phi}(f_{n})=\mathcal{L}_{\phi}(f_{n})-\mathcal{L}_{\phi}(f^{*}_{\phi})\leq a_{n}.$
  3. (A3)
    There exists a sequence $\{B_{n}\}_{n\in\mathbb{N}}$ with $B_{n}\gtrsim 1$ such that
    $\sup_{f\in\mathcal{F}_{n}}\|f\|_{L^{\infty}([0,1]^{d})}\leq B_{n}.$
  4. (A4)
    There exists a constant $\nu\in(0,1]$ such that, for any $f\in\mathcal{F}_{n}$,
    $\mathbb{E}_{{\boldsymbol{X}},Y}\left[(\phi(Yf({\boldsymbol{X}}))-\phi(Yf^{*}_{\phi}({\boldsymbol{X}})))^{2}\right]\leq c_{2}B_{n}^{2-\nu}\mathcal{R}_{\phi}(f)^{\nu},$
    where $c_{2}>0$ is a constant depending only on $\phi$ and the conditional class probability function $\eta$.
  5. (A5)
    There exist a sequence $\{\delta_{n}\}_{n\in\mathbb{N}}$ and a constant $c_{3}>0$ such that
    $\log\mathcal{N}(\delta_{n},\mathcal{F}_{n},\|\cdot\|_{L^{\infty}([0,1]^{d})})\leq c_{3}n\left(\frac{\delta_{n}}{B_{n}}\right)^{2-\nu}.$
Let $\epsilon_{n}\asymp a_{n}\lor\delta_{n}$ and $\widehat{f}_{\phi,n}$ be the empirical $\phi$-risk minimizer (4.2) over the function class $\mathcal{F}_{n}$. Then,
$\mathbb{P}(\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\geq\epsilon_{n})\lesssim\exp(-c_{4}n(\epsilon_{n}/B_{n})^{2-\nu}),$
for some constant $c_{4}>0$. In particular, if $n(\epsilon_{n}/B_{n})^{2-\nu}\gtrsim(\log n)^{1+r}$ for some $r>0$, then
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\right]\lesssim\epsilon_{n}.$
Lemma 4.3 provides a systematic way to derive convergence rates of the excess $\phi$-risk for general surrogate losses. For the hinge loss $\phi(t)=\max\{1-t,0\}$ and the logistic loss $\phi(t)=\log(1+e^{-t})$, the condition (A1) is satisfied with $c_{1}=1$. When the function class $\mathcal{F}_{n}$ is parameterized by a convolutional neural network, the covering number bound in the condition (A5) can be checked by using Theorem 2.5. Note that $a_{n}$ in the condition (A2) quantifies how well $f^{*}_{\phi}$ can be approximated by $\mathcal{F}_{n}$ in the $\phi$-loss.
For the hinge loss, we know that $f^{*}_{\phi}=\,{\rm sgn}\,(2\eta-1)$. Since $f^{*}_{\phi}$ is bounded, we can simply choose $B_{n}=1$ in the condition (A3). The variance bound in (A4) was established by [Wu and Zhou, 2005; Steinwart and Scovel, 2007, Lemma 6.1]. We summarize their result in the following lemma. It shows that we can choose $\nu=q/(q+1)$ in the condition (A4) of Lemma 4.3.
Lemma 4.4.
Assume the noise condition (4.3) holds for some $q\in[0,\infty]$. Let $\phi$ be the hinge loss. For any $f:[0,1]^{d}\to\mathbb{R}$ satisfying $\|f\|_{L^{\infty}([0,1]^{d})}\leq B$, it holds that
$\mathbb{E}_{{\boldsymbol{X}},Y}\left[(\phi(Yf({\boldsymbol{X}}))-\phi(Yf^{*}_{\phi}({\boldsymbol{X}})))^{2}\right]\leq c_{\eta,q}(B+1)^{\frac{q+2}{q+1}}\mathcal{R}_{\phi}(f)^{\frac{q}{q+1}},$
where $c_{\eta,q}>0$ is a constant depending on $\eta$ and $q$.
Under the assumption that $\eta\in\mathcal{H}^{\alpha}(R)$, we can use Theorem 2.1 to construct a CNN $h$ that approximates $12\eta-$ 1. We can then approximate $f^{*}_{\phi}=\,{\rm sgn}\,(2\eta-1)$ by a CNN of the form $g\circ h$, where $g$ is a piece-wise linear function that approximates the sign function. The approximation error can be estimated by using the expression (4.5). Under the noise condition (4.3), we proves that one can choose $a_{n}\asymp L_{n}^{-(q+1)\alpha/d}$ in the condition (A2). Furthermore, Theorem 2.5 shows that we can choose $\delta_{n}\asymp(L_{n}n^{-1}\log n)^{\frac{q+1}{q+2}}$ in (A5). The trade-off between $a_{n}$ and $\delta_{n}$ tells us how to choose the depth $L_{n}$ and gives the desired rate in Theorem 4.1.
For the logistic loss, the convergence rate in Theorem 4.2 can be derived in a similar manner. The following lemma shows that one can choose $\nu=1$ in the condition (A4) of Lemma 4.3.
Lemma 4.5.
Let $\phi$ be the logistic loss. For any $f:[0,1]^{d}\to\mathbb{R}$ satisfying $\|f\|_{L^{\infty}([0,1]^{d})}\leq B$ with $B\geq 2$, it holds that
$\mathbb{E}_{{\boldsymbol{X}},Y}\left[(\phi(Yf({\boldsymbol{X}}))-\phi(Yf^{*}_{\phi}({\boldsymbol{X}})))^{2}\right]\leq 3B\mathcal{R}_{\phi}(f).$
Proof.
Since $\eta({\boldsymbol{X}})=\mathbb{P}(Y=1|{\boldsymbol{X}})$,
$\mathbb{E}_{{\boldsymbol{X}},Y}\left[(\phi(Yf({\boldsymbol{X}}))-\phi(Yf^{*}_{\phi}({\boldsymbol{X}})))^{2}\right]$
$=$ $\mathbb{E}_{{\boldsymbol{X}},Y}\left[\log^{2}\left(\frac{1+e^{-Yf({\boldsymbol{X}})}}{1+e^{-Yf^{*}_{\phi}({\boldsymbol{X}})}}\right)\right]$
$=$ $\mathbb{E}_{\boldsymbol{X}}\left[\eta({\boldsymbol{X}})\log^{2}\left(\frac{1+e^{-f({\boldsymbol{X}})}}{1+e^{-f^{*}_{\phi}({\boldsymbol{X}})}}\right)+(1-\eta({\boldsymbol{X}}))\log^{2}\left(\frac{1+e^{f({\boldsymbol{X}})}}{1+e^{f^{*}_{\phi}({\boldsymbol{X}})}}\right)\right]$
$=$ $\mathbb{E}_{\boldsymbol{X}}\left[\eta({\boldsymbol{X}})\log^{2}\left(\frac{\eta({\boldsymbol{X}})}{\psi(f({\boldsymbol{X}}))}\right)+(1-\eta({\boldsymbol{X}}))\log^{2}\left(\frac{1-\eta({\boldsymbol{X}})}{1-\psi(f({\boldsymbol{X}}))}\right)\right],$
where $\psi$ is the logistic function defined by (4.7) and we use $f^{*}_{\phi}=\log(\frac{\eta}{1-\eta})$ in the last equality. Since $|f({\boldsymbol{X}})|\leq B$ with $B\geq 2$, we have $\psi(f({\boldsymbol{X}}))\in[\psi(-B),\psi(B)]=[\psi(-B),1-\psi(-B)]$ with $\psi(-B)\leq e^{-2}$. We can apply the following inequalities proven in Proposition A.3 in the Appendix:
$\eta\log^{2}\left(\frac{\eta}{\psi(f)}\right)$ $\leq 2\log(1+e^{B})\left(\eta\log\left(\frac{\eta}{\psi(f)}\right)-\eta+\psi(f)\right),$
$(1-\eta)\log^{2}\left(\frac{1-\eta}{1-\psi(f)}\right)$ $\leq 2\log(1+e^{B})\left((1-\eta)\log\left(\frac{1-\eta}{1-\psi(f)}\right)+\eta-\psi(f)\right),$
where we omit the variable ${\boldsymbol{X}}$. Thus,
$\mathbb{E}_{{\boldsymbol{X}},Y}\left[(\phi(Yf({\boldsymbol{X}}))-\phi(Yf^{*}_{\phi}({\boldsymbol{X}})))^{2}\right]$
$\leq$ $2\log(1+e^{B})\mathbb{E}\left[\eta\log\left(\frac{\eta}{\psi(f)}\right)+(1-\eta)\log\left(\frac{1-\eta}{1-\psi(f)}\right)\right]$
$\leq$ $3B\mathcal{R}_{\phi}(f),$
where we use the equality (4.8) and $\log(1+e^{B})\leq 3B/2$. ∎
Recall that, for the logistic loss, $f^{*}_{\phi}=\log(\frac{\eta}{1-\eta})$ and $\eta=\psi(f^{*}_{\phi})$, where $\psi$ is defined by (4.7). Since $\eta\in\mathcal{H}^{\alpha}(R)$, it can be approximated by a CNN $h$ by using Theorem 2.1. We further construct a CNN $f=g\circ h$, where $g$ is a function that approximates the mapping $t\mapsto\log t-\log(1-t)$, such that $\psi(f)$ approximates $\eta$ well. By equality (4.8), the excess $\phi$-risk $\mathcal{R}_{\phi}(f)=\mathbb{E}[\mathcal{D}_{KL}(\eta,\psi(f))]$ can be estimated by using the following lemma, which is a modification from [Bos and Schmidt-Hieber, 2022, Theorem 3.2].
Lemma 4.6.
Let $u\in(0,1/2)$ and suppose the SVB condition (4.10) holds for some $\beta\in[0,1]$. If a function $h:[0,1]^{d}\to[u,1-u]$ satisfies $\|h-\eta\|_{L^{\infty}([0,1]^{d})}\leq Cu$ for some constant $C>0$, then
$\mathbb{E}_{\boldsymbol{X}}[\mathcal{D}_{KL}(\eta({\boldsymbol{X}}),h({\boldsymbol{X}}))]\leq\begin{cases}\frac{2(2-\beta)C_{\beta}(C+1)^{2+\beta}}{1-\beta}u^{1+\beta},&\beta<1,\\ 2C_{1}(C+1)^{3}u^{2}\log(u^{-1}),&\beta=1.\end{cases}$
Using Lemma 4.6, we prove that one can choose $a_{n}\asymp L_{n}^{-(1+\beta)\alpha/d}\log n$ and $B_{n}\asymp\log n$ in conditions (A2) and (A3) of Lemma 4.3. Since $\nu=1$, Theorem 2.5 shows that we can choose $\delta_{n}\asymp L_{n}B_{n}n^{-1}\log n$ in (A5). The trade-off between $a_{n}$ and $\delta_{n}$ tells us how to choose the depth $L_{n}$ and gives the rate in Theorem 4.2.
5 Conclusion
In this paper, we have studied approximation and learning capacities of convolutional neural networks with one-side zero-padding and multiple channels. We have derived new approximation bounds for CNNs with norm constraint on the weights. To study the generalization performance of learning algorithms induced by these networks, we also proved new bounds for their covering number. Based on these results, we established rates of convergence for CNNs in nonparametric regression and classification problems. Many of the obtained convergence rates are known to be minimax optimal.
There is a restriction on the smoothness of the target functions in our results. We think this restriction is due to the proof techniques of our approximation bound (Theorem 2.1), rather than the architecture of CNNs. It may be possible to use the ideas of network constructions from related works, such as [Oono and Suzuki, 2019], to remove the restriction, which we leave as a future work.
Acknowledgments
The work described in this paper was partially supported by Discovery Project (DP240101919) of the Australian Research Council, InnoHK initiative, The Government of the HKSAR, Laboratory for AI-Powered Financial Technologies, the Research Grants Council of Hong Kong [Projects Nos. CityU 11315522, CityU 11303821] and National Natural Science Foundation of China [Project No. 12371103].
Appendix A Appendix
A.1 Basic properties of CNNs
In this section, we give some properties of the function class $\mathcal{CNN}(s,J,L,M)$, which will be useful for neural network construction.
Proposition A.1.
If $J\leq J^{\prime}$, $L\leq L^{\prime}$ and $M\leq M^{\prime}$, then $\mathcal{CNN}(s,J,L,M)\subseteq\mathcal{CNN}(s,J^{\prime},L^{\prime},M^{\prime})$.
Proof.
It is easy to check that $\mathcal{CNN}(s,J,L,M)\subseteq\mathcal{CNN}(s,J^{\prime},L,M^{\prime})$ by adding zero filters appropriately and using the definition of weight constraint (2.3). To prove $\mathcal{CNN}(s,J^{\prime},L,M^{\prime})\subseteq\mathcal{CNN}(s,J^{\prime},L^{\prime},M^{\prime})$, we observe that the convolution with the filter $(1,0,\dots,0)^{\intercal}\in\mathbb{R}^{s}$ is the identity map of $\mathbb{R}^{d}\to\mathbb{R}^{d}$. Hence, we can increase the depth of CNN by adding the identity map in the last layer. ∎
The next proposition shows that we can always rescale the parameters in CNNs so that the norms of filters in the hidden layers are at most one.
Proposition A.2 (Rescaling).
Every $f\in\mathcal{CNN}(s,J,L,M)$ can be parameterized in the form (2.1) such that $\|{\boldsymbol{w}}^{(L)}\|_{1}\leq M$ and $\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|\leq 1$ for all $\ell\in[0:L-1]$.
Proof.
The proof is essentially the same as [Jiao et al., 2023, Proposition 2.4]. We give the proof here for completeness. Note that $f\in\mathcal{CNN}(s,J,L,M)$ parameterized in the form (2.1) can be written inductively by
$f({\boldsymbol{x}})=\langle{\boldsymbol{w}}^{(L)},{\boldsymbol{f}}_{L}({\boldsymbol{x}})\rangle,\quad{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})=\sigma\left(\,{\rm Conv}\,_{{\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))\right),\quad{\boldsymbol{f}}_{0}({\boldsymbol{x}})={\boldsymbol{x}}.$
Let us denote $m_{\ell}:=\max\{\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|,1\}$ for all $\ell\in[0:L-1]$ and let $\widetilde{{\boldsymbol{w}}}_{(\ell)}={\boldsymbol{w}}_{\ell}/m_{\ell}$, $\widetilde{{\boldsymbol{b}}}_{\ell}={\boldsymbol{b}}_{\ell}/(\prod_{i=0}^{\ell}m_{i})$ and $\widetilde{{\boldsymbol{w}}}_{(L)}={\boldsymbol{w}}^{(L)}\prod_{i=0}^{L-1}m_{i}$. We consider the functions defined inductively by
$\widetilde{{\boldsymbol{f}}}_{\ell+1}({\boldsymbol{x}})=\sigma\left(\,{\rm Conv}\,_{\widetilde{{\boldsymbol{w}}}^{(\ell)},\widetilde{{\boldsymbol{b}}}^{(\ell)}}(\widetilde{{\boldsymbol{f}}}_{\ell}({\boldsymbol{x}}))\right),\quad\widetilde{{\boldsymbol{f}}}_{0}({\boldsymbol{x}})={\boldsymbol{x}}.$
It is easy to check that $\|\widetilde{{\boldsymbol{w}}}_{(L)}\|\leq M$ and
$\left\|(\widetilde{{\boldsymbol{w}}}^{(\ell)},\widetilde{{\boldsymbol{b}}}^{(\ell)})\right\|=\frac{1}{m_{\ell}}\left\|\left({\boldsymbol{w}}^{(\ell)},\frac{{\boldsymbol{b}}^{(\ell)}}{\prod_{i=0}^{\ell-1}m_{i}}\right)\right\|\leq\frac{1}{m_{\ell}}\left\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\right\|\leq 1,$
where the first inequality is due to $m_{i}\geq 1$.
Next, we show that ${\boldsymbol{f}}_{\ell}({\boldsymbol{x}})=\left(\prod_{i=0}^{\ell-1}m_{i}\right)\widetilde{{\boldsymbol{f}}}_{\ell}({\boldsymbol{x}})$ by induction. For $\ell=1$, by the absolute homogeneity of the ReLU function,
${\boldsymbol{f}}_{1}({\boldsymbol{x}})$ $=\sigma\left(\,{\rm Conv}\,_{{\boldsymbol{w}}^{(0)},{\boldsymbol{b}}^{(0)}}({\boldsymbol{x}})\right)$
$=m_{0}\sigma\left(\,{\rm Conv}\,_{\widetilde{{\boldsymbol{w}}}^{(0)},\widetilde{{\boldsymbol{b}}}^{(0)}}({\boldsymbol{x}})\right)=m_{0}\widetilde{{\boldsymbol{f}}}_{1}({\boldsymbol{x}}).$
Inductively, one can conclude that
${\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})$ $=\sigma\left(\,{\rm Conv}\,_{{\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))\right)$
$=\left(\prod_{i=0}^{\ell}m_{i}\right)\sigma\left(\,{\rm Conv}\,_{\widetilde{{\boldsymbol{w}}}^{(\ell)},\widetilde{{\boldsymbol{b}}}^{(\ell)}}\left(\frac{{\boldsymbol{f}}_{\ell}({\boldsymbol{x}})}{\prod_{i=0}^{\ell-1}m_{i}}\right)\right)$
$=\left(\prod_{i=0}^{\ell}m_{i}\right)\sigma\left(\,{\rm Conv}\,_{\widetilde{{\boldsymbol{w}}}^{(\ell)},\widetilde{{\boldsymbol{b}}}^{(\ell)}}\left(\widetilde{{\boldsymbol{f}}}_{\ell}({\boldsymbol{x}})\right)\right)$
$=\left(\prod_{i=0}^{\ell}m_{i}\right)\widetilde{{\boldsymbol{f}}}_{\ell+1}({\boldsymbol{x}}),$
where the third equality is due to induction. Therefore,
$f({\boldsymbol{x}})=\langle{\boldsymbol{w}}^{(L)},{\boldsymbol{f}}_{L}({\boldsymbol{x}})\rangle=\left\langle{\boldsymbol{w}}^{(L)},\left(\prod_{i=0}^{L-1}m_{i}\right)\widetilde{{\boldsymbol{f}}}_{L}({\boldsymbol{x}})\right\rangle=\langle\widetilde{{\boldsymbol{w}}}^{(L)},\widetilde{{\boldsymbol{f}}}_{L}({\boldsymbol{x}})\rangle,$
which means $f$ can be parameterized by $(\widetilde{{\boldsymbol{w}}}^{(0)},\widetilde{{\boldsymbol{b}}}^{(0)},\dots,\widetilde{{\boldsymbol{w}}}^{(L-1)},\widetilde{{\boldsymbol{b}}}^{(L-1)},\widetilde{{\boldsymbol{w}}}^{(L)})$ and we finish the proof. ∎
A.2 A useful inequality
Proposition A.3.
If $p\in[0,1]$ and $q\in[u,1]$ with $0<u\leq e^{-2}$, then
$p\log^{2}(p/q)\leq\log(u^{-2})(p\log(p/q)-p+q).$
Proof.
It is easy to check that the inequality holds for $p=0$, since $00\log^{2}0=0\log 0=$ 0. So, we only consider $p>0$. If we denote $t=q/p\geq u$, then the desired inequality is equivalent to
$t-1-\log t\geq\frac{\log^{2}t}{-2\log u}.$ (A.1)
It is easy to see that this inequality holds for $t=1$. For $t\in(1,\infty)$, it can be proven by letting $t=e^{s}$ with $s>0$ and using
$e^{s}-1-s\geq\frac{s^{2}}{2}\geq\frac{s^{2}}{-2\log u}.$
For $t\in[u,1)$, we consider the function
$f(t)=\frac{\log^{2}t}{t-1-\log t}.$
Since $t-1-\log t>0$, the inequality (A.1) is equivalent to $f(t)\leq-2\log u$. A direct calculation shows
$f^{\prime}(t)$ $=\frac{\log t}{(t-1-\log t)^{2}}\left(2-\frac{2}{t}-\log t-\frac{\log t}{t}\right)$
$=:\frac{\log t}{(t-1-\log t)^{2}}g(t).$
Since
$g^{\prime}(t)=\frac{1-t+\log t}{t^{2}}<0,$
we know that $g(t)>g(1)=0$ and hence $f^{\prime}(t)<0$ for $t\in[u,1)$. Therefore, $f$ is decreasing on $[u,1)$ and
$f(t)\leq f(u)=\frac{\log^{2}u}{u-1-\log u}\leq-2\log u,$
where we use $u-1\geq\log(e^{-1})\geq\frac{1}{2}\log u$ because $u\leq e^{-2}$. ∎
A.3 Proof of Theorem 2.1
We first construct a CNN to implement the function of the form ${\boldsymbol{x}}\mapsto c\sigma({\boldsymbol{a}}^{\intercal}{\boldsymbol{x}}+b)$.
Lemma A.4.
Let $s\in[2:d]$ and $L=\lceil\frac{d-1}{s-1}\rceil$. For any ${\boldsymbol{a}}\in\mathbb{R}^{d}$ and $b,c\in\mathbb{R}$, there exists $f\in\mathcal{CNN}(s,3,L,M)$ such that $f({\boldsymbol{x}})=c\sigma({\boldsymbol{a}}^{\intercal}{\boldsymbol{x}}+b)$ for ${\boldsymbol{x}}\in[0,1]^{d}$ and $M=3^{L-1}|c|(\|{\boldsymbol{a}}\|_{1}+|b|)$. Furthermore, the output weights ${\boldsymbol{w}}^{(L)}\in\mathbb{R}^{d\times 3}$ can be chosen to satisfy $w^{(L)}_{i,j}=0$ except for $i=j=1$.
Proof.
By the homogeneity of ReLU, we can assume that $\|{\boldsymbol{a}}\|_{1}+|b|=1$. Observe that convolution with the filter $(u_{1},\dots,u_{s})^{\intercal}\in\mathbb{R}^{s}$ computes the inner product with the first $s$ elements of the input signal $\sum_{i\in[s]}u_{i}x_{i}$. Convolution with the filter $(0,\dots,0,1)^{\intercal}\in\mathbb{R}^{s}$ is the “left-translation” by $s-1$. We construct the CNN parameterized in the form (2.1) as follows.
We define ${\boldsymbol{w}}^{(0)}\in\mathbb{R}^{s\times 3\times 1}$ and ${\boldsymbol{b}}^{(0)}\in\mathbb{R}^{3}$ by
$w^{(0)}_{:,1,1}=-w^{(0)}_{:,2,1}=\begin{pmatrix}a_{1}\\ \vdots\\ a_{s}\end{pmatrix},\quad w^{(0)}_{:,3,1}=\begin{pmatrix}0\\ \vdots\\ 0\\ 1\end{pmatrix},\quad{\boldsymbol{b}}^{(0)}=0.$
Then, the pre-activated output of the first layer, i.e. $\,{\rm Conv}\,_{{\boldsymbol{w}}^{(0)},{\boldsymbol{b}}^{(0)}}({\boldsymbol{x}})$, is
$\begin{pmatrix}\sum_{i\in[s]}a_{i}x_{i}&-\sum_{i\in[s]}a_{i}x_{i}&x_{s}\\ *&*&\vdots\\ \vdots&\vdots&x_{d}\\ *&*&*\end{pmatrix}\in\mathbb{R}^{d\times 3},$
where we use *** to denote some entries that we do not care. By using the equality $t=\sigma(t)-\sigma(-t)$, we are able to compute the partial inner product $\sum_{i\in[s]}a_{i}x_{i}$ after applying the ReLU activation. Since we assume ${\boldsymbol{x}}\in[0,1]^{d}$, the $x_{i}=\sigma(x_{i})$, $i\in[s:d]$, are stored in the output of the first layer. For $\ell\in[L-2]$, we define ${\boldsymbol{w}}^{(\ell)}\in\mathbb{R}^{s\times 3\times 3}$ and ${\boldsymbol{b}}^{(\ell)}\in\mathbb{R}^{3}$ by
$w^{(\ell)}_{:,1,:}=-w^{(\ell)}_{:,2,:}=\begin{pmatrix}1&-1&0\\ 0&0&a_{\ell(s-1)+2}\\ \vdots&\vdots&\vdots\\ 0&0&a_{(\ell+1)(s-1)+1}\end{pmatrix},\quad w^{(\ell)}_{:,3,:}=\begin{pmatrix}0&0&0\\ \vdots&\vdots&\vdots\\ 0&0&0\\ 0&0&1\end{pmatrix},\quad{\boldsymbol{b}}^{(\ell)}=0.$
Then, the pre-activated output of the $\ell+1$-th layer is
$\begin{pmatrix}\sum_{i\in[(\ell+1)(s-1)+1]}a_{i}x_{i}&-\sum_{i\in[(\ell+1)(s-1)+1]}a_{i}x_{i}&x_{(\ell+1)(s-1)+1}\\ *&*&\vdots\\ \vdots&\vdots&x_{d}\\ *&*&*\end{pmatrix}\in\mathbb{R}^{d\times 3}.$
For $\ell=L-1$, we let ${\boldsymbol{b}}^{(L-1)}=(b,0,0)^{\intercal}$. The filter ${\boldsymbol{w}}^{(L-1)}\in\mathbb{R}^{s\times 3\times 3}$ and the pre-activated output $o_{L}\in\mathbb{R}^{d\times 3}$ are given below
$w^{(L-1)}_{:,1,:}=\begin{pmatrix}1&-1&0\\ 0&0&a_{(L-1)(s-1)+2}\\ \vdots&\vdots&\vdots\\ \vdots&\vdots&a_{d}\\ 0&0&0\\ \end{pmatrix},\quad w^{(L-1)}_{:,2,:}=w^{(\ell)}_{:,3,:}=0,\quad o_{L}=\begin{pmatrix}{\boldsymbol{a}}^{\intercal}{\boldsymbol{x}}+b&0&0\\ *&0&0\\ \vdots&\vdots&\vdots\\ *&0&0\end{pmatrix}.$
Finally, for the output weights ${\boldsymbol{w}}^{(L)}\in\mathbb{R}^{d\times 3}$, we let $w^{(L)}_{1,1}=c$ and $w^{(L)}_{i,j}=0$ otherwise. Then, the output of the CNN is exactly $c\sigma({\boldsymbol{a}}^{\intercal}{\boldsymbol{x}}+b)$. It is easy to see that $\|({\boldsymbol{w}}^{(0)},{\boldsymbol{b}}^{(0)})\|\leq 1$, $\|{\boldsymbol{w}}^{(L)}\|_{1}=|c|$, $\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|\leq 3$ for $\ell\in[L-1]$. Hence, for this parameterization, $\kappa({\boldsymbol{\theta}})\leq 3^{L-1}|c|$. ∎
Next, we show that shallow neural networks defined by (2.4) can be parameterized by a CNN.
Lemma A.5.
Let $s\in[2:d]$ and $L_{0}=\lceil\frac{d-1}{s-1}\rceil$. Then, for any $f\in\mathcal{NN}(N,M)$, there exists $f_{\boldsymbol{\theta}}\in\mathcal{CNN}(s,6,NL_{0},3^{L_{0}+1}NM)$ such that $f_{\boldsymbol{\theta}}({\boldsymbol{x}})=f({\boldsymbol{x}})$ for all ${\boldsymbol{x}}\in[0,1]^{d}$.
Proof.
Given a parameter $R>0$, which will be chosen later, any function $f\in\mathcal{NN}(N,M)$ can be written as
$f({\boldsymbol{x}})=\frac{M}{R}\sum_{i=1}^{N}c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i}),$
where $\sum_{i=1}^{N}|c_{i}|(\|{\boldsymbol{a}}_{i}\|_{1}+|b_{i}|)\leq R$. By Lemma A.4, the function ${\boldsymbol{x}}\mapsto c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i})$ can be implemented by $\mathcal{CNN}(s,3,L_{0},3^{L_{0}-1}R)$, where $L_{0}=\lceil\frac{d-1}{s-1}\rceil$. We denote the corresponding parameters by $({\boldsymbol{w}}^{(0)}(i),{\boldsymbol{b}}^{(0)}(i),\dots,{\boldsymbol{w}}^{(L_{0}-1)}(i),{\boldsymbol{b}}^{(L_{0}-1)}(i),{\boldsymbol{w}}^{(L_{0})}(i))$, where $w^{(L_{0})}_{j,k}(i)=0$ except for $j=k=1$. By Proposition A.2, we can further assume that $|w^{(L_{0})}_{1,1}(i)|\leq 3^{L_{0}-1}R$ and $\|({\boldsymbol{w}}^{(\ell)}(i),{\boldsymbol{b}}^{(\ell)}(i))\|\leq 1$ for all $\ell\in[0:L_{0}-1]$.
In order to compute the summation $\sum c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i})$ in a sequential way, we use the 4th channel in the CNN to store the input ${\boldsymbol{x}}\in[0,1]^{d}$, and the 5th and 6th channels to store the partial summations of the positive part $\sum_{c_{i}>0}c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i})$ and the negative part $\sum_{c_{i}<0}-c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i})$, respectively. The filters ${\boldsymbol{w}}^{(\ell)}\in\mathbb{R}^{s\times 6\times 6}$ and bias ${\boldsymbol{b}}^{(\ell)}\in\mathbb{R}^{6}$ are defined as follows. For $\ell=0$,
$w^{(0)}_{:,1:3,1}={\boldsymbol{w}}^{(0)}(1),\quad w^{(0)}_{:,4,1}=(1,0,\dots,0)^{\intercal},\quad{\boldsymbol{b}}^{(0)}=({\boldsymbol{b}}^{(0)}(1)^{\intercal},0,0,0)^{\intercal}.$
Here and in the sequel, we use zero filters and biases when they are not specific defined. We always use the filter $w^{(\ell)}_{:,4,4}=w^{(\ell)}_{:,5,5}=w^{(\ell)}_{:,6,6}=(1,0,\dots,0)^{\intercal}$ to store the input and partial summations, except for the output layer. If $\ell=(i-1)L_{0}+j$ for some $i\in[N]$ and $j\in[L_{0}-1]$, then we define
$w^{(\ell)}_{:,1:3,1:3}={\boldsymbol{w}}^{(j)}(i),\quad{\boldsymbol{b}}^{(\ell)}=({\boldsymbol{b}}^{(j)}(i)^{\intercal},0,0,0)^{\intercal}.$
Recall that the only nonzero element in ${\boldsymbol{w}}^{(L_{0})}(i)$ is $w^{(L_{0})}_{1,1}(i)$. For $\ell=iL_{0}$ for some $i\in[N-1]$, we define
$w^{(\ell)}_{:,1:3,4}={\boldsymbol{w}}^{(0)}(i+1),\quad{\boldsymbol{b}}^{(\ell)}=({\boldsymbol{b}}^{(0)}(i+1)^{\intercal},0,0,0)^{\intercal},$
and
$w^{(\ell)}_{:,5,1}$ $=(w^{(L_{0})}_{1,1}(i),0,\cdots,0)^{\intercal},\quad\mbox{if }c_{i}>0,$
$w^{(\ell)}_{:,6,1}$ $=(-w^{(L_{0})}_{1,1}(i),0,\cdots,0)^{\intercal},\quad\mbox{if }c_{i}<0.$
Then, the activated output of the $iL_{0}$ layer is of the form
$\begin{pmatrix}c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i})/w^{(L_{0})}_{1,1}(i)&*&*&x_{1}&\sum\limits_{c_{j}>0,j<i}c_{j}\sigma({\boldsymbol{a}}_{j}^{\intercal}{\boldsymbol{x}}+b_{j})&\sum\limits_{c_{j}<0,j<i}-c_{j}\sigma({\boldsymbol{a}}_{j}^{\intercal}{\boldsymbol{x}}+b_{j})\\ *&*&*&\vdots&*&*\\ *&*&*&x_{d}&*&*\end{pmatrix},$
where *** denotes some entries that we do not care. It is easy to check that we correctly compute the partial summations $\sum_{c_{j}>0,j\leq i}c_{j}\sigma({\boldsymbol{a}}_{j}^{\intercal}{\boldsymbol{x}}+b_{j})$ and $-\sum_{c_{j}>0,j\leq i}c_{j}\sigma({\boldsymbol{a}}_{j}^{\intercal}{\boldsymbol{x}}+b_{j})$ in the $iL_{0}+1$ layer. Finally, for the output layer, i.e. $\ell=NL_{0}$, we define
${\boldsymbol{w}}^{(NL_{0})}=\frac{M}{R}\begin{pmatrix}w^{(L_{0})}_{1,1}(N)&0&0&0&1&-1\\ 0&0&0&0&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\end{pmatrix}.$
Then, the constructed CNN $f_{\boldsymbol{\theta}}({\boldsymbol{x}})$ implements the function $f({\boldsymbol{x}})$.
In our construction, $\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|=1+|w^{(L_{0})}_{1,1}(i)|\leq 1+3^{L_{0}-1}R$ if $\ell=iL_{0}$ for some $i\in[N-1]$ and $\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|\leq 1$ for other $\ell\in[0:NL_{0}-1]$, and $\|{\boldsymbol{w}}^{(NL_{0})}\|_{1}=MR^{-1}(2+|w^{(L_{0})}_{1,1}(N)|)\leq MR^{-1}(2+3^{L_{0}-1}R)$. Therefore, the norm constraint of the parameter is
$\kappa({\boldsymbol{\theta}})\leq\frac{M}{R}(2+3^{L_{0}-1}R)(1+3^{L_{0}-1}R)^{N-1}.$
If we choose $R=3^{1-L_{0}}N^{-1}$, then
$\kappa({\boldsymbol{\theta}})\leq\frac{2M}{R}(1+3^{L_{0}-1}R)^{N}\leq 3^{L_{0}+1}NM,$
where we use $(1+1/N)^{N}\leq e\leq 3$. ∎
We give a remark on the construction in the proof of Lemma A.5. This remark is useful for constructing CNNs to approximate functions of the form $g(f_{\boldsymbol{\theta}}({\boldsymbol{x}}))$ with $g:\mathbb{R}\to\mathbb{R}$.
Remark A.6.
We can replace the output layer (i.e. the parameters ${\boldsymbol{w}}^{(NL_{0})}$) of $f_{\boldsymbol{\theta}}$ by a convolutional layer with parameters $w^{(NL_{0})}_{:,1,:}=-w^{(NL_{0})}_{:,2,:}={\boldsymbol{w}}^{(NL_{0})}$ and $w^{(NL_{0})}_{:,i,:}=0$ for $i=3,4,5,6$. Then, the activated output of this CNN (without linear layer) is
$\begin{pmatrix}\sigma(f({\boldsymbol{x}}))&\sigma(-f({\boldsymbol{x}}))&0&0&0&0\\ *&*&\vdots&\vdots&\vdots&\vdots\\ *&*&0&0&0&0\end{pmatrix}.$ (A.2)
Thus, we can recover $f({\boldsymbol{x}})$ by using $\sigma(f({\boldsymbol{x}}))-\sigma(-f({\boldsymbol{x}}))$. The norm of this CNN can also be bounded by $3^{L_{0}+1}NM$.
Now, we can give a proof of Theorem 2.1 by using Lemma A.5.
Proof of Theorem 2.1.
Let $L_{0}=\lceil\frac{d-1}{s-1}\rceil$ and $N=\lfloor L/L_{0}\rfloor$. By Lemma A.5 and Proposition A.1, we have the inclusion $\mathcal{NN}(N,M_{0})\subseteq\mathcal{CNN}(s,6,L,M)$ for $M_{0}=3^{-L_{0}-1}N^{-1}M$. If $M\gtrsim L^{\frac{3d+3-2\alpha}{2d}}$, then $M_{0}\gtrsim L^{\frac{d+3-2\alpha}{2d}}$ and, by the approximation bound (2.5),
$\sup_{h\in\mathcal{H}^{\alpha}(1)}\inf_{f\in\mathcal{CNN}(s,6,L,M)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim N^{-\frac{\alpha}{d}}\lor M_{0}^{-\frac{2\alpha}{d+3-2\alpha}}\lesssim L^{-\frac{\alpha}{d}},$
which completes the proof. ∎
A.4 Proof of Theorem 4.1
Recall that, for the hing loss, $f^{*}_{\phi}=\,{\rm sgn}\,(2\eta-1)$. In order to estimate the approximation bound (A2) in Lemma 4.3, we need to construct a CNN to approximate $\,{\rm sgn}\,(2\eta-1)$. If $\eta$ is smooth, we can approximate $12\eta-$ 1 by a CNN $f$ using Theorem 2.1. The sign function can be approximated by a piece-wise linear function $g$. Hence, we need to construct a CNN to implement the composition $g\circ f$. The following lemma give such a construction. This lemma can be seen as an extension of Lemma A.5. Recall that we use $\mathcal{NN}(N,M)$ to denote the function class of shallow neural networks (2.4). We will use the notation $\mathcal{N}\mathcal{N}_{d}(N,M)$ to emphasize that the input dimension is $d$.
Lemma A.7.
Let $s\in[2:d]$, $L_{0}=\lceil\frac{d-1}{s-1}\rceil$ and $f\in\mathcal{N}\mathcal{N}_{d}(N,M)$. If $g\in\mathcal{N}\mathcal{N}_{1}(K,M_{0})$, then there exists $f_{\boldsymbol{\theta}}\in\mathcal{CNN}(s,6,NL_{0}+K+1,36\cdot 3^{L_{0}}NMKM_{0})$ such that $f_{\boldsymbol{\theta}}({\boldsymbol{x}})=g(f({\boldsymbol{x}}))$ for all ${\boldsymbol{x}}\in[0,1]^{d}$.
Proof.
Using the construction in the proof of Lemma A.5 and Remark A.6, we can construct $NL_{0}$ convolutional layers such that the activated output is given by (A.2) and the norm of these layers is at most $3^{L_{0}+1}NM$. Thus, we only need to implement the one-dimensional function $g$ by a $d$-dimensional CNN.
Without loss of generality, for any $R>0$, we can normalize $g$ such that
$g(t)=\frac{2M_{0}}{R}\sum_{k=1}^{K}c_{k}\sigma(a_{k}t+b_{k}),\quad|a_{k}|+|b_{k}|=\frac{1}{2},\sum_{k=1}^{K}|c_{k}|\leq R.$
The construction is similar to Lemma A.5. We use 2nd and 3rd channels channels to store the input information $\sigma(f({\boldsymbol{x}}))$ and $\sigma(-f({\boldsymbol{x}}))$. The 4th and 5th channels are used to store the partial summations of the positive part $\sum_{c_{k}>0}c_{k}\sigma(a_{k}f({\boldsymbol{x}})+b_{k})$ and the negative part $\sum_{c_{k}<0}-c_{k}\sigma(a_{k}f({\boldsymbol{x}})+b_{k})$, respectively. These can be done in a way similar to the proof of Lemma A.5. So, we only give the bias and the filter for the first channel: ${\boldsymbol{b}}^{(NL_{0}+k)}=(b_{k},0,0,0,0)^{\intercal}$,
$w^{(NL_{0}+k)}_{:,1,:}=\begin{pmatrix}0&a_{k}&-a_{k}&0&0\\ 0&0&0&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\end{pmatrix}.$
Then, the activated output of the $NL_{0}+k$-th layer is
$\begin{pmatrix}\sigma(a_{k}f({\boldsymbol{x}})+b_{k})&\sigma(f({\boldsymbol{x}}))&\sigma(-f({\boldsymbol{x}}))&\sum\limits_{c_{i}>0,i<k}c_{i}\sigma(a_{i}f({\boldsymbol{x}})+b_{i})&\sum\limits_{c_{i}<0,i<k}-c_{i}\sigma(a_{i}f({\boldsymbol{x}})+b_{i})\\ *&*&*&*&*\end{pmatrix},$
where, as before, *** denotes some entries that we do not care. Finally, the output layer is given by ${\boldsymbol{w}}^{(L)}\in\mathbb{R}^{d\times 5}$ with $L=NL_{0}+K+1$ and nonzero entries ${\boldsymbol{w}}^{(L)}_{1,1}=2M_{0}c_{K}/R$ and ${\boldsymbol{w}}^{(L)}_{1,4}=-{\boldsymbol{w}}^{(L)}_{1,5}=2M_{0}/R$.
The norm of the construed CNN is
$\kappa({\boldsymbol{\theta}})$ $\leq 3^{L_{0}+1}NM\cdot\frac{2M_{0}}{R}(2+|c_{K}|)\prod_{k=1}^{K-1}(1+|c_{k}|)$
$\leq 12\cdot 3^{L_{0}}NMM_{0}\frac{(1+R)^{K}}{R}$
$\leq 36\cdot 3^{L_{0}}NMKM_{0},$
where we choose $R=1/K$ and use $(1+1/K)^{K}\leq e\leq 3$ in the last inequality. ∎
We are ready to prove Theorem 4.1 by using Lemma 4.3.
Proof of Theorem 4.1.
Since $\eta\in\mathcal{H}^{\alpha}(R)$ by assumption, we have $2\eta-1\in\mathcal{H}^{\alpha}(2R+1)$. For a given $u_{n}\in(0,1)$, by the approximation bound (2.5), there exists $h\in\mathcal{NN}(\widetilde{N}_{n},\widetilde{M}_{n})$ with $\widetilde{N}_{n}\asymp u_{n}^{-d/\alpha}$ and $\widetilde{M}_{n}\gtrsim u_{n}^{-(d+3-2\alpha)/(2\alpha)}$ such that
$\|2\eta-1-h\|_{L^{\infty}([0,1]^{d})}\leq u_{n}.$ (A.3)
Recall that $f^{*}_{\phi}=\,{\rm sgn}\,(2\eta-1)$ for the hing loss. We define
$g(t)=\begin{cases}1\quad&t\geq u_{n},\\ t/u_{n}\quad&|t|<u_{n},\\ -1\quad&t\leq-u_{n},\end{cases}$
which can also be written as
$g(t)=u_{n}^{-1}\sigma(t+u_{n})-u_{n}^{-1}\sigma(t-u_{n})-\sigma(0t+1).$
Since $J\geq 6$ and $2u_{n}^{-1}(1+u_{n})+1\lesssim u_{n}^{-1}$, by Lemma A.7, there exists $f\in\mathcal{CNN}(s,J,L_{n},M_{n})$ with $L_{n}\asymp\widetilde{N}_{n}\asymp u_{n}^{-d/\alpha}$ and $M_{n}\asymp\widetilde{N}_{n}\widetilde{M}_{n}u_{n}^{-1}\gtrsim u_{n}^{-(3d+3)/(2\alpha)}$ such that $f=g\circ h$ on $[0,1]^{d}$. Notice that $\pi_{1}f=f$ by construction, which implies $f\in\mathcal{F}_{n}$.
Let us denote $\Omega_{n}:=\{{\boldsymbol{x}}\in[0,1]^{d}:|2\eta({\boldsymbol{x}})-1|\geq 2u_{n}\}$, then $|h({\boldsymbol{x}})|\geq u_{n}$ on $\Omega_{n}$ by (A.3). As a consequence, $f^{*}_{\phi}({\boldsymbol{x}})=\,{\rm sgn}\,(2\eta({\boldsymbol{x}})-1)=g(h({\boldsymbol{x}}))=f({\boldsymbol{x}})$ for any ${\boldsymbol{x}}\in\Omega_{n}$. Using equality (4.5), we get
$\mathcal{R}_{\phi}(f)$ $=\mathbb{E}[|f-f^{*}_{\phi}||2\eta-1|]$
[…]
$\leq 4u_{n}\mathbb{P}_{\boldsymbol{X}}(|2\eta({\boldsymbol{X}})-1|\leq 2u_{n})\lesssim u_{n}^{q+1}.$
Thus, we have shown that one can choose $a_{n}\asymp u_{n}^{q+1}$ and $B_{n}=1$ in conditions (A2) and (A3) of Lemma 4.3. By Lemma 4.4, we can choose $\nu=q/(q+1)$ in condition (A4) of Lemma 4.3. To select $\delta_{n}$ in condition (A5) of Lemma 4.3, we apply Theorem 2.5 to estimate the entropy of $\mathcal{F}_{n}$ (note that $\pi_{1}$ does not increase the entropy):
$\log\mathcal{N}(\delta_{n},\mathcal{F}_{n},\|\cdot\|_{L^{\infty}([0,1]^{d})})$ $\lesssim L_{n}\log(L_{n}M_{n}^{2}/\delta_{n})$
$\lesssim u_{n}^{-d/\alpha}\log(nu_{n}^{-1}\delta_{n}^{-1}),$
where we use $M_{n}\lesssim\,{\rm Poly}\,(n)$ in the last inequality. This bound shows that we can choose $\delta_{n}\asymp(n^{-1}u_{n}^{-d/\alpha}\log n)^{\frac{q+1}{q+2}}$ under the assumption that $u_{n}^{-1}\lesssim\,{\rm Poly}\,(n)$. Finally, if we set
$u_{n}\asymp\left(\frac{\log^{2}n}{n}\right)^{\frac{\alpha}{(q+2)\alpha+d}},$
then $\epsilon_{n}\asymp a_{n}\lor\delta_{n}\asymp a_{n}=u_{n}^{q+1}$ satisfies the condition $n(\epsilon_{n}/B_{n})^{2-\nu}\gtrsim\log^{2}n$ in Lemma 4.3, and hence we have
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\right]\lesssim\epsilon_{n}\lesssim\left(\frac{\log^{2}n}{n}\right)^{\frac{(q+1)\alpha}{(q+2)\alpha+d}},$
which completes the proof. ∎
A.5 Proof of Theorem 4.2
Recall that we use $\mathcal{N}\mathcal{N}_{1}(N,M)$ to denote the function class of shallow neural networks (2.4) with one dimensional input.
Lemma A.8.
For any integer $N\geq 3$, there exists $g\in\mathcal{N}\mathcal{N}_{1}(2N,6N)$ such that the following conditions hold
  1. (1)
    $|g(t)|\leq\log N$ for all $t\in\mathbb{R}$.
  2. (2)
    $g(t)=-\log N$ for $t\leq 0$, and $g(t)=\log N$ for $t\geq 1$.
  3. (3)
    For any $t\in[0,1]$, $|\psi(g(t))-t|\leq 3N^{-1}$, where $\psi$ is the logistic function (4.7).
Proof.
Let us first construct a function $h$ that approximates the logarithm function on $[0,1]$. For convenience, we denote $t_{i}=i/N$ for $i\in[0:N]$. We let $h(t)=\log t_{1}=-\log N$ for $t\leq t_{1}$ and $h(t)=0$ for $t\geq t_{N}=1$. In the interval $[t_{1},t_{N}]$, we let $h$ be the continuous piecewise linear function with $N-1$ pieces such that $h(t_{i})=\log t_{i}$ for $i\in[1:N]$. The function $h$ can be written as
$h(t)=-\log N+k_{1}\sigma(t-t_{1})-k_{N-1}\sigma(t-t_{N})+\sum_{i=2}^{N-1}(k_{i}-k_{i-1})\sigma(t-t_{i}),$
where $k_{i}=N(\log t_{i+1}-\log t_{i})\leq t_{i}^{-1}$ is the slope of $h$ on the interval $(t_{i},t_{i+1})$. Let us consider the approximation error $f(t)=\log t-h(t)$. By construction, $f(t)\geq 0$ on $[t_{1},t_{N}]$ and $f(t_{i})=0$ for $i\in[1:N]$. Observe that $|f^{\prime}(t)-f^{\prime}(x)|=|t^{-1}-x^{-1}|\leq t_{i}^{-2}|t-x|$ for $t,x\in[t_{i},t_{i+1}]$, $i\in[1:N-1]$. It is well known that such a Lipschitz gradient property implies the following inequality [Nesterov, 2018, Lemma 1.2.3]
$|f(x)-f(t)-f^{\prime}(t)(x-t)|\leq\frac{1}{2t_{i}^{2}}|x-t|^{2},\quad x,t\in[t_{i},t_{i+1}].$ (A.4)
If the function $f(t)$ attains its maximal value on $[t_{i},t_{i+1}]$ at some point $t_{i}^{*}\in(t_{i},t_{i+1})$, then $f^{\prime}(t_{i}^{*})=0$ and hence, by choosing $x=t_{i}$ and $t=t_{i}^{*}$ in (A.4), we get
$f(t_{i}^{*})\leq\frac{1}{2t_{i}^{2}}|t_{i}-t_{i}^{*}|^{2}\leq\frac{1}{2t_{i}^{2}N^{2}}.$
As a consequence, for $t\in[t_{i},t_{i+1}]$ with $i\in[1:N-1]$, we have
$|e^{h(t)}-t|$ $=t(1-e^{-f(t)})\leq tf(t)\leq\frac{t_{i}+\frac{1}{N}}{2t_{i}^{2}N^{2}}$
$=\frac{1}{2t_{i}N^{2}}+\frac{1}{2t_{i}^{2}N^{3}}\leq\frac{1}{N},$
where we use $t_{i}N=i\geq 1$ in the last inequality. Observe that, for $t\in[0,t_{1}]$, we have $|e^{h(t)}-t|=|t_{1}-t|\leq N^{-1}$. We conclude that $|e^{h(t)}-t|\leq N^{-1}$ holds for any $t\in[0,1]$.
Next, we define $g(t)=h(t)-h(1-t)$, then $g\in\mathcal{N}\mathcal{N}_{1}(2N,M)$, where
$M$ $\leq 3k_{1}+3k_{N-1}+3\sum_{i=2}^{N-1}(k_{i-1}-k_{i})$
$=6k_{1}\leq 6t_{1}^{-1}=6N.$
It is easy to check that conditions (1) and (2) hold. For $t\in[0,1]$,
$|\psi(g(t))-t|$ $=\left|\frac{e^{h(t)}}{e^{h(t)}+e^{h(1-t)}}-t\right|$
$\leq\left|e^{h(t)}-t\right|+\left|\frac{e^{h(t)}}{e^{h(t)}+e^{h(1-t)}}-e^{h(t)}\right|$
$=\left|e^{h(t)}-t\right|+\frac{e^{h(t)}}{e^{h(t)}+e^{h(1-t)}}\left|t-e^{h(t)}+1-t-e^{h(1-t)}\right|$
$\leq 3N^{-1},$
which proves condition (3). ∎
Proof of Theorem 4.2.
As in the proof of Theorem 4.1, for a given $u_{n}\in(0,1/3)$, there exists $h\in\mathcal{NN}(\widetilde{N}_{n},\widetilde{M}_{n})$ with $\widetilde{N}_{n}\asymp u_{n}^{-d/\alpha}$ and $\widetilde{M}_{n}\gtrsim u_{n}^{-(d+3-2\alpha)/(2\alpha)}$ such that
$\|\eta-h\|_{L^{\infty}([0,1]^{d})}\leq u_{n}.$
Let $g\in\mathcal{N}\mathcal{N}_{1}(2\lceil u_{n}^{-1}\rceil,6\lceil u_{n}^{-1}\rceil)$ be a function that satisfies Lemma A.8 with $N=\lceil u_{n}^{-1}\rceil$. Thus, $|g(t)|\leq\log\lceil u_{n}^{-1}\rceil$ for all $t\in\mathbb{R}$, and
$|\psi(g(t))-t|\leq 3u_{n}\quad t\in[0,1].$
Since $J\geq 6$, by Lemma A.7, there exists $f\in\mathcal{CNN}(s,J,L_{n},M_{n})$ with $L_{n}\asymp\widetilde{N}_{n}+\lceil u_{n}^{-1}\rceil\asymp u_{n}^{-d/\alpha}$ and $M_{n}\asymp\widetilde{N}_{n}\widetilde{M}_{n}\lceil u_{n}^{-1}\rceil^{2}\gtrsim u_{n}^{-(3d+3+2\alpha)/(2\alpha)}$ such that $f=g\circ h$ on $[0,1]^{d}$. If we let $B_{n}=\log\lceil u_{n}^{-1}\rceil$, then $\pi_{B_{n}}f=f$ by construction, which implies $f\in\mathcal{F}_{n}$.
Define the function $\widetilde{h}:[0,1]^{d}\to[0,1]$ by
$\widetilde{h}({\boldsymbol{x}})=\begin{cases}0,&h({\boldsymbol{x}})<0,\\ h({\boldsymbol{x}}),&h({\boldsymbol{x}})\in[0,1],\\ 1,&h({\boldsymbol{x}})>1.\end{cases}$
By the second condition in Lemma A.8 for the function $g$, we have $f=g\circ h=g\circ\widetilde{h}$. Since $\eta({\boldsymbol{x}})\in[0,1]$ for any ${\boldsymbol{x}}\in[0,1]^{d}$, we also have $|\widetilde{h}({\boldsymbol{x}})-\eta({\boldsymbol{x}})|\leq|h({\boldsymbol{x}})-\eta({\boldsymbol{x}})|\leq u_{n}$. Therefore,
$|\psi(f({\boldsymbol{x}}))-\eta({\boldsymbol{x}})|$ $\leq|\psi(g(\widetilde{h}({\boldsymbol{x}})))-\widetilde{h}({\boldsymbol{x}})|+|\widetilde{h}({\boldsymbol{x}})-\eta({\boldsymbol{x}})|$
$\leq 3u_{n}+u_{n}=4u_{n}.$
Since $\psi(f({\boldsymbol{x}}))\in[\psi(-B_{n}),\psi(B_{n})]\subseteq[u_{n}/2,1-u_{n}/2]$, Lemma 4.6 implies that
$\mathcal{R}_{\phi}(f)=\mathbb{E}[\mathcal{D}_{KL}(\eta,\psi(f))]\lesssim u_{n}^{1+\beta}\log u_{n}^{-1}.$
Thus, we have shown that one can choose $a_{n}\asymp u_{n}^{1+\beta}\log u_{n}^{-1}$ and $B_{n}=\log\lceil u_{n}^{-1}\rceil$ in conditions (A2) and (A3) of Lemma 4.3. By Lemma 4.5, we can choose $\nu=1$ in condition (A4) of Lemma 4.3. Using Theorem 2.5, we can estimate the entropy of $\mathcal{F}_{n}$ as
$\log\mathcal{N}(\delta_{n},\mathcal{F}_{n},\|\cdot\|_{L^{\infty}([0,1]^{d})})$ $\lesssim L_{n}\log(L_{n}M_{n}^{2}/\delta_{n})$
$\lesssim u_{n}^{-d/\alpha}\log(nu_{n}^{-1}\delta_{n}^{-1}).$
where we use $M_{n}\lesssim\,{\rm Poly}\,(n)$ in the last inequality. This bound shows that we can choose $\delta_{n}\asymp n^{-1}u_{n}^{-d/\alpha}B_{n}\log n$ in condition (A5) of Lemma 4.3, under the assumption that $u_{n}^{-1}\lesssim\,{\rm Poly}\,(n)$. Finally, if we set
$u_{n}\asymp\left(\frac{\log n}{n}\right)^{\frac{\alpha}{(1+\beta)\alpha+d}},$
then $\epsilon_{n}\asymp a_{n}\lor\delta_{n}\asymp a_{n}\asymp\delta_{n}$ satisfies the condition $n\epsilon_{n}/B_{n}\gtrsim\log^{2}n$ in Lemma 4.3, and hence we have
$\mathbb{E}\left[\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\right]\lesssim\epsilon_{n}\lesssim\left(\frac{\log n}{n}\right)^{\frac{(1+\beta)\alpha}{(1+\beta)\alpha+d}}\log n,$
which completes the proof. ∎
A.6 Proof of Lemma 4.6
Using the inequality $\log t\leq t-1$ for $t>0$, we have for any ${\boldsymbol{x}}\in[0,1]^{d}$,
$\mathcal{D}_{KL}(\eta({\boldsymbol{x}}),h({\boldsymbol{x}}))$ $\leq\eta({\boldsymbol{x}})\left(\frac{\eta({\boldsymbol{x}})}{h({\boldsymbol{x}})}-1\right)+(1-\eta({\boldsymbol{x}}))\left(\frac{1-\eta({\boldsymbol{x}})}{1-h({\boldsymbol{x}})}-1\right)$
$=\frac{(\eta({\boldsymbol{x}})-h({\boldsymbol{x}}))^{2}}{h({\boldsymbol{x}})(1-h({\boldsymbol{x}}))}\leq\frac{C^{2}u^{2}}{h({\boldsymbol{x}})}+\frac{C^{2}u^{2}}{1-h({\boldsymbol{x}})}.$ (A.5)
By assumption, we have $h({\boldsymbol{x}})\geq u$ and $h({\boldsymbol{x}})\geq\eta({\boldsymbol{x}})-Cu$. Notice that, if $\eta({\boldsymbol{x}})-Cu\geq u$, then
$\eta({\boldsymbol{x}})-Cu\geq\eta({\boldsymbol{x}})-\frac{C\eta({\boldsymbol{x}})}{C+1}=\frac{\eta({\boldsymbol{x}})}{C+1}.$
Therefore,
$\frac{1}{h({\boldsymbol{x}})}\leq\frac{1}{u}\mathbbm{1}_{\{\eta({\boldsymbol{x}})<(C+1)u\}}+\frac{C+1}{\eta({\boldsymbol{x}})}\mathbbm{1}_{\{\eta({\boldsymbol{x}})\geq(C+1)u\}}.$
By the SVB condition (4.10), we get
$\mathbb{E}_{\boldsymbol{X}}[h({\boldsymbol{X}})^{-1}]\leq C_{\beta}(C+1)^{\beta}u^{\beta-1}+(C+1)I_{u},$
where
[…]
If $\beta<1$, then by the SVB condition (4.10),
[…]
If $\beta=1$, the SVB condition (4.10) implies that $C_{\beta}\geq 1$ and $\mathbb{P}_{\boldsymbol{X}}(\eta({\boldsymbol{X}})\leq t^{-1})\leq\min\{1,C_{\beta}t^{-1}\}$, which leads to
[…]
Thus, we have given a bound for $\mathbb{E}_{\boldsymbol{X}}[h({\boldsymbol{X}})^{-1}]$. Similarly, one can show that the same bound holds for $\mathbb{E}_{\boldsymbol{X}}[(1-h({\boldsymbol{X}}))^{-1}]$. Combining these bounds with (A.5) finishes the proof.