### v3

# On the rates of convergence for learning with convolutional neural networks

Abstract
We study approximation and learning capacities of convolutional neural networks (CNNs) with one-side zero-padding and multiple channels. Our first result proves a new approximation bound for CNNs with certain constraint on the weights. Our second result gives new analysis on the covering number of feed-forward neural networks with CNNs as special cases. The analysis carefully takes into account the size of the weights and hence gives better bounds than the existing literature in some situations. Using these two results, we are able to derive rates of convergence for estimators based on CNNs in many learning problems. In particular, we establish minimax optimal convergence rates of the least squares based on CNNs for learning smooth functions in the nonparametric regression setting. For binary classification, we derive convergence rates for CNN classifiers with hinge loss and logistic loss. It is also shown that the obtained rates for classification are minimax optimal in some common settings.
keywords:
convolutional neural network, convergence rate, regression, classification, approximation, covering number
{MSCcodes}
68T07, 41A25, 62G08
1 Introduction
Deep leaning has made remarkable successes in many applications and research fields such as image classification, speech recognition, natural language processing and scientific computing [17, 9]. This breakthrough of deep learning also motivated many theoretical researches on understanding and explaining the empirical successes of deep neural networks from various perspectives. In particular, recent studies have established optimal approximations of smooth function classes by fully connected neural networks [38, 39, 27, 20]. It has also been shown that these networks can achieve minimax optimal rates of convergence in many learning problems, including nonparametric regression [25, 15] and classification [13].
Our main interests in this paper are the approximation and learning properties of convolutional neural networks (CNNs), which are widely used in image classification and related applications [16]. Recently, substantial progress has been made in the theoretical study of CNNs. It has been shown that CNNs are universal for approximation [43] and universally consistent for nonparametric regression [18]. Approximation bounds and representational advantages of CNNs have been proven in several works [24, 42, 7, 22]. Furthermore, rates of convergence of estimators based on CNNs were established for nonparametric regression [44, 37] and classification [14, 19, 8]. However, in contrast with the minimax optimal learning rates for fully connected neural networks [25, 15, 13], many of these results for CNNs are not optimal.
In this paper, we take a step to close this gap by providing new analysis on the approximation and learning capacities of CNNs. Specifically, we prove new bounds for the approximation of smooth functions by CNNs with certain constraint on the weights. We also derive bounds for the covering numbers of these networks. Using the obtained bounds, we are able to establish convergence rates for CNNs in many learning problems. These rates are known to be minimax optimal in several settings. We summarize our contributions in the following.
  1. (1)
    We prove the rates $\mathcal{O}(L^{-\alpha/d})$ for the approximation of smooth functions with smoothness $\alpha<(d+3)/2$ by CNNs, where $d$ is the dimension and $L$ is the depth of CNNs. The main advantage of our result is that we have an explicit control on the network weights (through $\kappa(\theta)$ defined by (3) below). It is proven that one can choose $\kappa(\theta)\leq M$ with $M=\mathcal{O}(L^{\frac{3d+3-2\alpha}{2d}})$ to ensure that the approximation rate $\mathcal{O}(L^{-\alpha/d})$ holds.
  2. (2)
    We provide a new framework to estimate the covering numbers of feed-forward neural networks. An application of this result gives the bound $\mathcal{O}(L\log(LM/\epsilon))$ for the $\epsilon$-covering number of CNNs with depth $L$ and weight constraint $\kappa(\theta)\leq M$. When $M$ grows at most polynomially on $L$, our bound is better than the general bound $\mathcal{O}(L^{2}\log(L/\epsilon))$ in the literature.
  3. (3)
    For regression, we establish the minimax optimal rate for the least squares regression with CNNs, when the regression function is smooth.
  4. (4)
    For binary classification, we establish rates of convergence for CNN classifiers with hinge loss and logistic loss, under the Tsybakov noise condition (15). For the hinge loss, the obtained rate for the excess classification risk is minimax optimal. For the logistic loss, the obtained rate may not be optimal for the excess classification risk. But it is optimal for the excess logistic risk, at least in some situations.
The remainder of this paper is organized as follows. In Section 2, we describe the architecture of convolutional neural networks used in this paper, and derive bounds for the approximation capacity and covering number of these networks. Sections 3 and 4 study the nonparametric regression and classification problems, respectively. We give convergence rates of the excess risk for CNNs in these two sections. Section 5 concludes this paper with a discussion on future studies. Omitted proofs are given in Supplementary Materials.
1.1 Notations
[…]
2 Convolutional neural networks
Let us first define convolutional neural networks used in this paper. Let ${\boldsymbol{w}}=(w_{1},\dots,w_{s})^{\intercal}\in\mathbb{R}^{s}$ be a filter with filter size $s\in[d]$. We define the convolution matrix $T_{\boldsymbol{w}}$ on $\mathbb{R}^{d}$ by
$T_{\boldsymbol{w}}:=\begin{pmatrix}w_{1}&\cdots&w_{s-1}&w_{s}\\ &\ddots&\ddots&\ddots&\ddots\\ &&w_{1}&\cdots&w_{s-1}&w_{s}\\ &&&w_{1}&\cdots&w_{s-1}\\ &&&&\ddots&\vdots\\ &&&&&w_{1}\end{pmatrix}\in\mathbb{R}^{d\times d}$
This convolution matrix corresponds to the one-sided padding and stride-one convolution by the filter ${\boldsymbol{w}}$. It is essentially the same as the convolution used in [24] (up to a matrix transpose). But it is different from the convolution matrix used in [42, 43, 7, 22, 8, 37], which is of dimension $(d+s)\times d$, rather than $d\times d$. So, in their setting, the network width increases after every application of the convolution with only one channel, while the network width remains the same with more channels in our setting. We define convolutional layers as follows. Let $s,J,J^{\prime}\in\mathbb{N}$ be a filter size, input channel size, and output channel size. For a filter ${\boldsymbol{w}}=(w_{i,j^{\prime},j})_{i\in[s],j^{\prime}\in[J^{\prime}],j\in[J]}\in\mathbb{R}^{s\times J^{\prime}\times J}$ and a bias vector ${\boldsymbol{b}}=(b_{1},\dots,b_{J^{\prime}})^{\intercal}\in\mathbb{R}^{J^{\prime}}$, we define the convolutional layer as an operator $\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}:\mathbb{R}^{d\times J}\to\mathbb{R}^{d\times J^{\prime}}$ by
$(\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}}))_{:,j^{\prime}}:=\sum_{j=1}^{J}T_{w_{:,j^{\prime},j}}x_{:,j}+b_{j^{\prime}},\quad{\boldsymbol{x}}=(x_{i,j})_{i\in[d],j\in[J]}\in\mathbb{R}^{d\times J},$
where we use “$+b_{j^{\prime}}$” to denote the vector addition “$+(b_{j^{\prime}},\dots,b_{j^{\prime}})^{\intercal}$” when there is no confusion. Next, we define convolutional neural networks. Let $s\in[d]$ and $J,L\in\mathbb{N}$ be the filter size, channel size and depth. We denote by $\mathcal{CNN}(s,J,L)$ the set of functions $f_{\boldsymbol{\theta}}$ that can be parameterized by ${\boldsymbol{\theta}}=({\boldsymbol{w}}^{(0)},{\boldsymbol{b}}^{(0)},\dots,{\boldsymbol{w}}^{(L-1)},{\boldsymbol{b}}^{(L-1)},{\boldsymbol{w}}^{(L)})$ in the following form
(1) $f_{\boldsymbol{\theta}}({\boldsymbol{x}}):=\left\langle{\boldsymbol{w}}^{(L)},\sigma\circ\,{\rm Conv}\,_{{\boldsymbol{w}}^{(L-1)},{\boldsymbol{b}}^{(L-1)}}\circ\cdots\circ\sigma\circ\,{\rm Conv}\,_{{\boldsymbol{w}}^{(0)},{\boldsymbol{b}}^{(0)}}({\boldsymbol{x}})\right\rangle,\quad{\boldsymbol{x}}\in[0,1]^{d},$
where ${\boldsymbol{w}}^{(0)}\in\mathbb{R}^{s\times J\times 1},{\boldsymbol{b}}^{(0)}\in\mathbb{R}^{J},{\boldsymbol{w}}^{(L)}\in\mathbb{R}^{d\times J},{\boldsymbol{w}}^{(\ell)}\in\mathbb{R}^{s\times J\times J},{\boldsymbol{b}}^{(\ell)}\in\mathbb{R}^{J}$ for $\ell\in[L-1]$, and the activation $\sigma(t)=t\lor 0$ is the ReLU activation function. Note that we have assumed the channel sizes in each layers are the same, because we can always increase the channel sizes by adding appropriate zero filters and biases. For convenience, we will often view ${\boldsymbol{w}}^{(0)}\in\mathbb{R}^{s\times J\times J}$ by adding zeros to the filter and the input. The number of parameters in the network is $(sJ+1)JL+(d+s-sJ)J\lesssim J^{2}L$, which grows linearly on the depth $L$.
In order to control the complexity of convolutional neural networks, we introduce the following norm for the pair $({\boldsymbol{w}},{\boldsymbol{b}})\in\mathbb{R}^{s\times J\times J}\times\mathbb{R}^{J}$
$\|({\boldsymbol{w}},{\boldsymbol{b}})\|:=\max_{j^{\prime}\in[J]}\left(\left\|w_{:,j^{\prime},:}\right\|_{1}+|b_{j^{\prime}}|\right).$
Note that $\|({\boldsymbol{w}},{\boldsymbol{b}})\|$ quantifies the size of the affine transform $\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}$:
$\left\|\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}})\right\|_{\infty}$ $\leq\max_{j^{\prime}\in[J]}\left(\sum_{j=1}^{J}\left\|T_{w_{:,j^{\prime},j}}x_{:,j}\right\|_{\infty}+|b_{j^{\prime}}|\right)$
(2) $\leq\|({\boldsymbol{w}},{\boldsymbol{b}})\|(\|{\boldsymbol{x}}\|_{\infty}\lor 1).$
Following the idea of [12], we define a constraint on the weights as follows
(3) $\kappa({\boldsymbol{\theta}}):=\|{\boldsymbol{w}}^{(L)}\|_{1}\prod_{\ell=0}^{L-1}\left(\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|\lor 1\right).$
For any $M\geq 0$, we denote the function class consisting of wights constrained CNNs by
$\mathcal{CNN}(s,J,L,M):=\left\{f_{\boldsymbol{\theta}}\in\mathcal{CNN}(s,J,L):\kappa({\boldsymbol{\theta}})\leq M\right\}.$
Several properties of this function class are summarized in Section LABEL:sec:_property of Supplementary Materials. In Sections 2.1 and 2.2, we study the approximation capacity and covering number of $\mathcal{CNN}(s,J,L,M)$. These results are used in Sections 3 and 4 to study the convergence rates of CNNs on the nonparametric regression and classification problems.
2.1 Approximation
We consider the capacity of CNNs for approximating smooth functions. Given a smoothness index $\alpha>0$, we write $\alpha=r+\beta$ where $r\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ and $\beta\in(0,1]$. Let $C^{r,\beta}(\mathbb{R}^{d})$ be the Hölder space with the norm
$\|f\|_{C^{r,\beta}(\mathbb{R}^{d})}:=\max\left\{\|f\|_{C^{r}(\mathbb{R}^{d})},\max_{\|{\boldsymbol{s}}\|_{1}=r}|\partial^{\boldsymbol{s}}f|_{C^{0,\beta}(\mathbb{R}^{d})}\right\},$
where ${\boldsymbol{s}}=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index and
$\|f\|_{C^{r}(\mathbb{R}^{d})}$ $:=\max_{\|{\boldsymbol{s}}\|_{1}\leq r}\|\partial^{\boldsymbol{s}}f\|_{L^{\infty}(\mathbb{R}^{d})},$
$|f|_{C^{0,\beta}(\mathbb{R}^{d})}$ $:=\sup_{{\boldsymbol{x}}\neq{\boldsymbol{y}}\in\mathbb{R}^{d}}\frac{|f({\boldsymbol{x}})-f({\boldsymbol{y}})|}{\|{\boldsymbol{x}}-{\boldsymbol{y}}\|_{2}^{\beta}}.$
Here, we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm since we only consider continuous functions. We write $C^{r,\beta}([0,1]^{d})$ for the Banach space of all restrictions to $[0,1]^{d}$ of functions in $C^{r,\beta}(\mathbb{R}^{d})$. The norm is given by $\|f\|_{C^{r,\beta}([0,1]^{d})}=\inf\{\|g\|_{C^{r,\beta}(\mathbb{R}^{d})}:g\in C^{r,\beta}(\mathbb{R}^{d})\mbox{ and }g=f\mbox{ on }[0,1]^{d}\}$. For convenience, we will denote the ball of $C^{r,\beta}([0,1]^{d})$ with radius $R>0$ by
$\mathcal{H}^{\alpha}(R):=\left\{f\in C^{r,\beta}([0,1]^{d}):\|f\|_{C^{r,\beta}([0,1]^{d})}\leq R\right\}.$
Note that, for $\alpha=1$, $\mathcal{H}^{1}(R)$ is a class of Lipschitz continuous functions.
Our first result estimates the error of approximating Hölder functions by CNNs.
Theorem 2.1.
Let $0<\alpha<(d+3)/2$ and $s\in[2:d]$. If $L\geq\lceil\frac{d-1}{s-1}\rceil$ and $M\gtrsim L^{\frac{3d+3-2\alpha}{2d}}$, then
$\sup_{h\in\mathcal{H}^{\alpha}(1)}\inf_{f\in\mathcal{CNN}(s,6,L,M)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim L^{-\frac{\alpha}{d}}.$
The approximation rate $\mathcal{O}(L^{-\alpha/d})$ is slightly better than $\mathcal{O}((L/\log L)^{-\alpha/d})$ in [24, Corollary 4] and [19, Theorem 1] for ResNet-type CNNs. Furthermore, the results of [24, 19] requires that the depth of residual blocks grows with the approximation error, while our result does not need any residual blocks. Our approximation rate is the same as the result of [37], which used slightly different CNNs, and the rate in [8], which considered the approximation of smooth functions on spheres. There is another recent paper [26] proving the so-called super-convergence rate $\mathcal{O}((L/\log L)^{-2\alpha/d})$ for CNNs by combining the super-convergence rate for fully-connected networks [20, 11] and the result of [42], which showed that fully-connected networks can be implemented by CNNs. Note that the network architecture in [26] is also different to ours because they need downsampling layers in order to apply [42, Theorem 2]. We summarize and compare the network architectures and approximation results of these papers in Table 1.
Table 1: A comparison of network architectures and approximation results for CNNs in recent works. The rate of approximation by CNNs with depth $L$ is given for target functions with smoothness $\alpha$ and input dimension $d$. “Residual”, “Downsampling” and “FC” mean that residual blocks, downsampling layers and fully-connected layers are used, respectively.
Network architecture Weight constraint Target function Approximation rate
[24] CNN+Residual maximum magnitude Hölder $(L/\log L)^{-\alpha/d}$
[19] CNN+Residual maximum magnitude Besov $(L/\log L)^{-\alpha/d}$
[8] CNN+Downsampling+FC None Sobolev on sphere $L^{-\alpha/d}$
[26] CNN+Downsampling None Sobolev $(L/\log L)^{-2\alpha/d}$
[37] CNN None Hölder, $\alpha<(d+3)/2$ $L^{-\alpha/d}$
Ours CNN $\kappa(\theta)$ defined by (3) Hölder, $\alpha<(d+3)/2$ $L^{-\alpha/d}$
Approximation results for neural networks can be divided into two categories according to whether the network weights are constrained. When there is no weight constraint, one can derive super-convergence rate by using the bit extraction technique [3, 20, 28] and estimate the complexity of the network using VC-dimension [3, 10, 15]. However, if the magnitudes of the weights are constrained, it seems that one can only get the slow rate $\mathcal{O}(L^{-\alpha/d})$. In this case, we can directly estimate the covering number of the network [8, 24, 25]. Comparing with existing results, which often bound the maximum magnitude of the weights, the main advantage of Theorem 2.1 is that we provide an explicitly bound on the weight constraint $\kappa(\theta)\leq M$, which leads to an optimal estimate of the covering number as shown by Theorem 2.8 below.
It is difficult to directly construct CNNs to approximate smooth functions. As mentioned above, most existing works derive approximation rates for CNNs by using the idea that smooth functions are well approximated by fully-connected neural networks and one can construct CNNs to implement fully-connected neural networks [24, 43, 42, 26]. Our proof of Theorem 2.1 is also based on this idea. The main technical difference from previous works is that we apply the approximation bound for shallow neural networks proven in [37]. To be concrete, let us denote the function class of shallow neural networks by
(4) $\mathcal{NN}(N,M):=\left\{f({\boldsymbol{x}})=\sum_{i=1}^{N}c_{i}\sigma({\boldsymbol{a}}_{i}^{\intercal}{\boldsymbol{x}}+b_{i}):\sum_{i=1}^{N}|c_{i}|(\|{\boldsymbol{a}}_{i}\|_{1}+|b_{i}|)\leq M\right\}.$
It was shown by [37, Corollary 2.4] that, if $\alpha<(d+3)/2$, then
(5) $\sup_{h\in\mathcal{H}^{\alpha}(1)}\inf_{f\in\mathcal{NN}(N,M)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim N^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+3-2\alpha}}.$
In order to apply the bound (5), we first construct a CNN to implement the function of the form ${\boldsymbol{x}}\mapsto c\sigma({\boldsymbol{a}}^{\intercal}{\boldsymbol{x}}+b)$. Different from previous works on CNNs [43, 42], we give an explicit estimate on the size of the weights.
Lemma 2.2.
Let $s\in[2:d]$ and $L=\lceil\frac{d-1}{s-1}\rceil$. For any ${\boldsymbol{a}}\in\mathbb{R}^{d}$ and $b,c\in\mathbb{R}$, there exists $f\in\mathcal{CNN}(s,3,L,M)$ such that $f({\boldsymbol{x}})=c\sigma({\boldsymbol{a}}^{\intercal}{\boldsymbol{x}}+b)$ for ${\boldsymbol{x}}\in[0,1]^{d}$ and $M=3^{L-1}|c|(\|{\boldsymbol{a}}\|_{1}+|b|)$. Furthermore, the output weights ${\boldsymbol{w}}^{(L)}\in\mathbb{R}^{d\times 3}$ can be chosen to satisfy $w^{(L)}_{i,j}=0$ except for $i=j=1$.
Using this result, we further show that shallow neural networks defined by (4) can be parameterized by a CNN in the following lemma. Theorem 2.1 is a direct consequence of the approximation bound (5) and this lemma. The detailed proofs are given in Supplementary Material.
Lemma 2.3.
Let $s\in[2:d]$ and $L_{0}=\lceil\frac{d-1}{s-1}\rceil$. Then, for any $f\in\mathcal{NN}(N,M)$, there exists $f_{\boldsymbol{\theta}}\in\mathcal{CNN}(s,6,NL_{0},3^{L_{0}+1}NM)$ such that $f_{\boldsymbol{\theta}}({\boldsymbol{x}})=f({\boldsymbol{x}})$ for all ${\boldsymbol{x}}\in[0,1]^{d}$.
Note that it is possible to extend the approximation bound (5) to Sobolev spaces with smoothness $\alpha\leq(d+3)/2$ by using the Radon transform as done in the recent paper [23]. Consequently, one can also generalize Theorem 2.1 and hence the statistical learning bounds in Sections 3 and 4 to these spaces by using the same argument. The restriction on the smoothness $\alpha<(d+3)/2$ is of course due to the use of the approximation bound (5). For high smoothness $\alpha>(d+3)/2$, one can also derive approximation bounds for CNNs by using the results of [37], as discussed in the following remark.
Remark 2.4.
If $\alpha>(d+3)/2$, it was shown by [37, Theorem 2.1] that $\mathcal{H}^{\alpha}(1)\subseteq\mathcal{F}_{\sigma}(M)$ for some constant $M>0$, where
[…]
Here, $\mathbb{S}^{d}$ is the unit sphere of $\mathbb{R}^{d+1}$ and $\|\mu\|=|\mu|(\mathbb{S}^{d})$ is the total variation of the measure $\mu$. $\mathcal{F}_{\sigma}(M)$ is a ball with radius $M$ of the variation space corresponding to shallow ReLU neural networks studied in many recent papers, such as [2, 6, 30, 31, 29]. The function class $\mathcal{F}_{\sigma}(M)$ can be viewed as output functions of an infinitely wide neural network. It is the limit of $\mathcal{NN}(N,M)$ as the number of neurons $N\to\infty$ [37, Proposition 2.2]. The recent work [29] showed that
$\sup_{h\in\mathcal{F}_{\sigma}(1)}\inf_{f\in\mathcal{NN}(N,1)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim N^{-\frac{d+3}{2d}}.$
Combining this bound with Lemma 2.3, we can obtain
(7) $\sup_{h\in\mathcal{F}_{\sigma}(1)}\inf_{f\in\mathcal{CNN}(s,6,L,M)}\|h-f\|_{L^{\infty}([0,1]^{d})}\lesssim L^{-\frac{d+3}{2d}},$
for $M\gtrsim L$. This approximation bound can be used to study machine learning problems with smoothness assumption $\alpha>(d+3)/2$, see Remark 3.2 for example. However, the bound (7) is not optimal for CNNs and high smoothness. In order to get a better approximation rate, a possible way is to first derive a bound similar to (5) for high smoothness and deep networks with proper constraint on the weights (see [12] for instance), and then show that these networks can be implemented by CNNs. We leave this for future study.
2.2 Covering number
In statistical learning theory, we often estimate generalization error of learning algorithms by certain complexities of models. The complexity we use in this paper is the covering number (or metric entropy) defined in the following.
Definition 2.5 (Covering number and entropy).
Let $\rho$ be a metric on a metric space $\mathcal{M}$ and $\mathcal{F}\subseteq\mathcal{M}$. For $\epsilon>0$, a set $\mathcal{S}\subseteq\mathcal{M}$ is called an $\epsilon$-cover (or $\epsilon$-net) of $\mathcal{F}$ if for any $x\in\mathcal{F}$, there exists $y\in\mathcal{S}$ such that $\rho(x,y)\leq\epsilon$. The $\epsilon$-covering number of $\mathcal{F}$ is defined by
$\mathcal{N}(\epsilon,\mathcal{F},\rho):=\min\{|\mathcal{S}|:\mathcal{S}\mbox{ is an $\epsilon$-cover of }\mathcal{F}\},$
where $|\mathcal{S}|$ is the cardinality of the set $\mathcal{S}$. The logarithm of the covering number $\log\mathcal{N}(\epsilon,\mathcal{F},\rho)$ is called (metric) entropy.
[…]
We consider neural networks of the following form
(8) ${\boldsymbol{f}}_{0}({\boldsymbol{x}})$ $={\boldsymbol{x}}\in[0,1]^{d},$
${\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})$ $=\sigma({\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))),\quad\ell\in[0:L-1],$
$f_{\boldsymbol{\theta}}({\boldsymbol{x}})$ $={\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{L}}({\boldsymbol{f}}_{L}({\boldsymbol{x}})),$
where ${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}:\mathbb{R}^{d_{\ell}}\to\mathbb{R}^{d_{\ell+1}}$ is an affine map parameterized by a vector ${\boldsymbol{\theta}}_{\ell}\in\mathbb{R}^{N_{\ell}}$ with $d_{0}=d$, $d_{L+1}=1$ and the vector of parameters ${\boldsymbol{\theta}}:=({\boldsymbol{\theta}}_{0}^{\intercal},\dots,{\boldsymbol{\theta}}_{L}^{\intercal})^{\intercal}\in\mathbb{R}^{N}$. Here, we use $N=\sum_{\ell=0}^{L}N_{\ell}$ to denote the number of parameters in the network. Note that we have restricted the input of the networks to $[0,1]^{d}$ for convenience. We assume that the parameterization satisfies the following conditions: for any ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in\mathbb{R}^{d_{\ell}}$ and $\ell\in[0:L]$,
(9) $\|{\boldsymbol{\theta}}\|_{\infty}$ $\leq B,$
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})\|_{\infty}$ $\leq\gamma_{\ell}(\|{\boldsymbol{x}}\|_{\infty}\lor 1),$
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}}^{\prime})\|_{\infty}$ $\leq\gamma_{\ell}\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|_{\infty},$
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}}({\boldsymbol{x}})\|_{\infty}$ $\leq\lambda_{\ell}\|{\boldsymbol{\theta}}_{\ell}-{\boldsymbol{\theta}}_{\ell}^{\prime}\|_{\infty}(\|{\boldsymbol{x}}\|_{\infty}\lor 1),$
where ${\boldsymbol{\theta}}^{\prime}$ denotes any parameters satisfying $\|{\boldsymbol{\theta}}^{\prime}\|_{\infty}\leq B$.
Note that, if the affine map have the following matrix form
${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{x}})={\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}}{\boldsymbol{x}}+{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}}=({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})\begin{pmatrix}{\boldsymbol{x}}\\ 1\end{pmatrix},$
then we can choose $\gamma_{\ell}$ to be the matrix operator norm (induced by $\|\cdot\|_{\infty}$)
$\gamma_{\ell}=\left\|({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})\right\|_{l^{\infty}\to l^{\infty}},$
and choose $\lambda_{\ell}$ to be the Lipschitz constant of the parameterization
$\left\|({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})-({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}})\right\|_{l^{\infty}\to l^{\infty}}\leq\lambda_{\ell}\|{\boldsymbol{\theta}}_{\ell}-{\boldsymbol{\theta}}_{\ell}^{\prime}\|_{\infty}.$
Recall that the matrix operator norm $\left\|({\boldsymbol{A}},{\boldsymbol{b}})\right\|_{l^{\infty}\to l^{\infty}}$ is the maximal 111-norm of rows of the matrix $({\boldsymbol{A}},{\boldsymbol{b}})$. In our constructions, $({\boldsymbol{A}}_{{\boldsymbol{\theta}}_{\ell}},{\boldsymbol{b}}_{{\boldsymbol{\theta}}_{\ell}})$ is linear on the parameter ${\boldsymbol{\theta}}_{\ell}$, which implies that we can choose $\lambda_{\ell}\lesssim d_{\ell}+1$.
The next lemma estimates the covering numbers of the neural networks described above.
Lemma 2.6.
Let $\mathcal{F}$ be the class of functions $f_{\boldsymbol{\theta}}$ that can be parameterized in the form (8), where the parameterization satisfies (9) with $\lambda_{\ell}\geq 0$ and $\gamma_{\ell}\geq 1$ for $\ell=[0:L]$. Then, the $\epsilon$-covering number of $\mathcal{F}$ in the $L^{\infty}([0,1]^{d})$ norm satisfies
$\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{L^{\infty}([0,1]^{d})})\leq(C_{L}B/\epsilon)^{N},$
where $N$ is the number of parameters and $C_{L}$ can be computed inductively by
$C_{0}=\lambda_{0},\quad C_{\ell+1}=\gamma_{\ell+1}C_{\ell}+\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}.$
In particular,
$C_{L}\leq\left(\sum_{j=0}^{L}\lambda_{j}\right)\prod_{i=0}^{L}\gamma_{i}.$
Proof 2.7.
For any ${\boldsymbol{\theta}},{\boldsymbol{\theta}}^{\prime}\in[-B,B]^{N}$ with $\|{\boldsymbol{\theta}}-{\boldsymbol{\theta}}^{\prime}\|_{\infty}\leq\epsilon$, we claim that, for any ${\boldsymbol{x}}\in[0,1]^{d}$ and $\ell\in[0:L]$,
$\|{\boldsymbol{f}}_{\ell}({\boldsymbol{x}})\|_{\infty}\leq\prod_{i=-1}^{\ell-1}\gamma_{i},$
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}}({\boldsymbol{f}}_{\ell}^{\prime}({\boldsymbol{x}}))\|_{\infty}\leq C_{\ell}\epsilon,$
$C_{\ell}\leq\left(\sum_{j=0}^{\ell}\lambda_{j}\right)\prod_{i=0}^{\ell}\gamma_{i},$
where we set $\gamma_{-1}=1$ and ${\boldsymbol{f}}_{\ell}^{\prime}$ denotes the function in (8) parameterized by ${\boldsymbol{\theta}}^{\prime}$. Thus, any $\epsilon$-cover of $[-B,B]^{N}$ gives a $C_{L}\epsilon$-cover of $\mathcal{F}$ in the $L^{\infty}([0,1]^{d})$ norm. Since the $\epsilon$-covering number of $[-B,B]^{N}$ is at most $(B/\epsilon)^{N}$, we get the desire bound for $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{L^{\infty}([0,1]^{d})})$.
We prove the claim by induction on $\ell\in[0:L]$. The claim is trivial for $\ell=0$ by definition. Assume that the claim is true for some $0\leq\ell<L$, we are going to prove it for $\ell+1$. By induction hypothesis,
$\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})\|_{\infty}\leq\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))\|_{\infty}\leq\gamma_{\ell}(\|{\boldsymbol{f}}_{\ell}({\boldsymbol{x}})\|_{\infty}\lor 1)\leq\prod_{i=0}^{\ell}\gamma_{i},$
where we used $\gamma_{i}\geq 1$ in the last inequality. By the Lipschitz continuity of ReLU,
$\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})-{\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}})\|_{\infty}\leq\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}({\boldsymbol{f}}_{\ell}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}^{\prime}}({\boldsymbol{f}}_{\ell}^{\prime}({\boldsymbol{x}}))\|_{\infty}\leq C_{\ell}\epsilon.$
Therefore,
$\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}}))\|_{\infty}$
$\leq$ $\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))\|_{\infty}+\|{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}}))-{\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell+1}^{\prime}}({\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}}))\|_{\infty}$
$\leq$ $\lambda_{\ell+1}\epsilon(\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})\|_{\infty}\lor 1)+\gamma_{\ell+1}\|{\boldsymbol{f}}_{\ell+1}({\boldsymbol{x}})-{\boldsymbol{f}}_{\ell+1}^{\prime}({\boldsymbol{x}})\|_{\infty}$
$\leq$ $\left(\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}+\gamma_{\ell+1}C_{\ell}\right)\epsilon=C_{\ell+1}\epsilon.$
Finally, by induction hypothesis and $\gamma_{\ell+1}\geq 1$,
$C_{\ell+1}$ $=\gamma_{\ell+1}C_{\ell}+\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}$
$\leq\left(\sum_{j=0}^{\ell}\lambda_{j}\right)\prod_{i=0}^{\ell+1}\gamma_{i}+\lambda_{\ell+1}\prod_{i=0}^{\ell}\gamma_{i}$
$\leq\left(\sum_{j=0}^{\ell+1}\lambda_{j}\right)\prod_{i=0}^{\ell+1}\gamma_{i},$
which completes the proof.
Now, we apply Lemma 2.6 to the convolutional neural network $\mathcal{CNN}(s,J,L,M)$. In this case, we have ${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{\ell}}=\,{\rm Conv}\,_{{\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)}}$ for $\ell\in[0:L-1]$ and ${\boldsymbol{\varphi}}_{{\boldsymbol{\theta}}_{L}}(\cdot)=\langle{\boldsymbol{w}}^{(L)},\cdot\rangle$. By Proposition LABEL:rescaling in Supplementary Materials, we can assume $\|{\boldsymbol{w}}^{(L)}\|_{1}\leq M$ and $\|({\boldsymbol{w}}^{(\ell)},{\boldsymbol{b}}^{(\ell)})\|\leq 1$ for all $\ell\in[0:L-1]$, which implies $B=M\lor 1$. Using the inequality (2) and
$\left\|\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}})-\,{\rm Conv}\,_{{\boldsymbol{w}},{\boldsymbol{b}}}({\boldsymbol{x}}^{\prime})\right\|_{\infty}$
$\leq$ $\max_{j^{\prime}\in[J]}\left(\sum_{j=1}^{J}\left\|T_{w_{:,j^{\prime},j}}x_{:,j}-T_{w_{:,j^{\prime},j}}x_{:,j}^{\prime}\right\|_{\infty}\right)$
$\leq$ $\|({\boldsymbol{w}},{\boldsymbol{b}})\|\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|_{\infty},$
we can set $\gamma_{\ell}=1$ and $\gamma_{L}=M$. It is easy to see that we can choose $\lambda_{\ell}=sJ+1$ and $\lambda_{L}=dJ$. Consequently,
$C_{L}\leq\left(\sum_{j=0}^{L}\lambda_{j}\right)\prod_{i=0}^{L}\gamma_{i}=(dJ+sJL+L)M\leq 3dJLM,$
where we use $s\leq d$ in the last inequality. We summarize the result in the next theorem for future reference.
Theorem 2.8.
Let $s,J,L\in\mathbb{N}$ and $M\geq 1$. The entropy of $\mathcal{CNN}(s,J,L,M)$ satisfies
$\log\mathcal{N}(\epsilon,\mathcal{CNN}(s,J,L,M),\|\cdot\|_{L^{\infty}([0,1]^{d})})\leq N\log(3dJLM^{2}/\epsilon),$
where $N=(sJ+1)JL+(d+s-sJ)J$ is the number of parameters in the network.
In the analysis of neural networks, many papers, such as [25, 8], simply assume that the parameters in the networks are bounded. In this case, the entropy is often bounded as $\mathcal{O}(NL\log(N/\epsilon))$, where $N$ is the number of parameters and $L$ is the depth. For convolutional neural networks with bounded width, we have $N\asymp L$ and hence the entropy would be $\mathcal{O}(L^{2}\log(L/\epsilon))$. For comparison, Theorem 2.8 gives a bound $\mathcal{O}(L\log(LM/\epsilon))$. If one only assumes that the parameters are bounded by $B$, then $M\lesssim B^{L}$ and our bound is consistent with the previous bound. However, if the weight constraint $M$ grows at most polynomially on $L$, then we get a better bound $\mathcal{O}(L\log(L/\epsilon))$ on the entropy. This improvement is essential to obtain optimal rates for many learning algorithms that we discuss in next two sections.
3 Regression
In this section, we consider the classical nonparametric regression problem. Assume that $({\boldsymbol{X}},Y)$ is a $[0,1]^{d}\times\mathbb{R}$-valued random vector satisfying $\mathbb{E}[Y^{2}]<\infty$. Let us denote the marginal distribution of ${\boldsymbol{X}}$ by $\mu$ and the regression function by
$h({\boldsymbol{x}}):=\mathbb{E}[Y|{\boldsymbol{X}}={\boldsymbol{x}}].$
Suppose we are given a data set of $n$ samples $\mathcal{D}_{n}=\{({\boldsymbol{X}}_{i},Y_{i})\}_{i=1}^{n}$, which are independent and have the same distribution as the random vector $({\boldsymbol{X}},Y)$. The goal of nonparametric regression problem is to construct an estimator $\widehat{f}_{n}$, based on $\mathcal{D}_{n}$, to reconstruct the regression function $h$. The estimation performance is evaluated by the $L^{2}$-error
$\|\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}=\mathbb{E}_{\boldsymbol{X}}\left[(\widehat{f}_{n}({\boldsymbol{X}})-h({\boldsymbol{X}}))^{2}\right].$
One of the popular algorithms to solve the regression problem is the empirical least squares
(10) $\widehat{f}_{n}\in\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}(f({\boldsymbol{X}}_{i})-Y_{i})^{2},$
where $\mathcal{F}_{n}$ is a prescribed hypothesis class. For simplicity, we assume here and in the sequel that the minimum above indeed exists. We are interested in the case that the function class $\mathcal{F}_{n}$ is parameterized by a CNN. In order to study the convergence rate of $\widehat{f}_{n}\to h$ as $n\to\infty$, we will assume that $h\in\mathcal{H}^{\alpha}(R)$ for some constant $R>0$ and make the following assumption on the distribution of $({\boldsymbol{X}},Y)$: there exists a constant $c>0$ such that
(11) $\mathbb{E}\left[\exp(cY^{2})\right]<\infty.$
In statistical analysis of learning algorithms, we often require that the hypothesis class is uniformly bounded. We define the truncation operator $\pi_{B}$ with level $B>0$ for real-valued functions $f$ as
(12) $\pi_{B}f({\boldsymbol{x}})=\begin{cases}B\quad&f({\boldsymbol{x}})>B,\\ f({\boldsymbol{x}})\quad&|f({\boldsymbol{x}})|\leq B,\\ -B\quad&f({\boldsymbol{x}})<-B.\end{cases}$
Note that the truncation operator can be implemented by a CNN (see Lemma LABEL:cnn_composition for example). Since we assume that the regression function $h$ is bounded, truncating the output of the estimator $\widehat{f}_{n}$ appropriately dose not increase the estimation error. The following theorem provides convergence rates for least squares estimators based on CNNs.
Theorem 3.1.
Assume that the condition (11) holds and the regression function $h\in\mathcal{H}^{\alpha}(R)$ for some $0<\alpha<(d+3)/2$ and $R>0$. Let $\widehat{f}_{n}$ be the estimator defined by (10) with $\mathcal{F}_{n}=\mathcal{CNN}(s,J,L_{n},M_{n})$, where $s\in[2:d]$, $J\geq 6$ and
$L_{n}\asymp\left(\frac{n}{\log^{3}n}\right)^{\frac{d}{2\alpha+d}},\quad\left(\frac{n}{\log^{3}n}\right)^{\frac{3d+3-2\alpha}{4\alpha+2d}}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n).$
If $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$, then
$\mathbb{E}_{\mathcal{D}_{n}}\left[\|\pi_{B_{n}}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]\lesssim\left(\frac{\log^{3}n}{n}\right)^{\frac{2\alpha}{2\alpha+d}}.$
It is well-known that the rate $n^{-\frac{2\alpha}{2\alpha+d}}$ is minimax optimal for learning functions in $\mathcal{H}^{\alpha}(R)$ [33]:
$\inf_{\widehat{f}_{n}}\sup_{h\in\mathcal{H}^{\alpha}(R)}\mathbb{E}_{\mathcal{D}_{n}}\left[\|\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]\gtrsim n^{-\frac{2\alpha}{2\alpha+d}},$
where the infimum is taken over all estimators based on the training data $\mathcal{D}_{n}$. Recent works have established the minimax rates (up to logarithm factors) for least squares estimators using fully-connected neural networks [25, 15, 36]. For convolutional neural networks, [24] proved the optimal rates for ResNet-type CNNs, under the requirement that the depth of the residual blocks grows with the sample size $n$, or the residual blocks are suitably masked. Theorem 3.1 removes the requirements on the residual blocks for low smoothness $\alpha<(d+3)/2$.
Remark 3.2.
As we noted in Remark 2.4, if $\alpha>(d+3)/2$, then $\mathcal{H}^{\alpha}(1)\subseteq\mathcal{F}_{\sigma}(R)$ for some constant $R>0$. When the regression function $h\in\mathcal{F}_{\sigma}(R)$, we can use the approximation bound (7) to show that
$\mathbb{E}_{\mathcal{D}_{n}}\left[\|\pi_{B_{n}}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\right]\lesssim\left(\frac{\log^{3}n}{n}\right)^{\frac{d+3}{2d+3}},$
if we choose $L_{n}\asymp(n/\log^{3}n)^{d/(2d+3)}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n)$. This rate is minimax optimal (up to logarithm factors) for the function class $\mathcal{F}_{\sigma}(R)$ [37]. For comparison, [37] only established the sub-optimal rate $\mathcal{O}(n^{-\frac{d+3}{3d+3}}\log^{4}n)$ for CNNs. Our result is also better than the recent analysis of CNNs in [44], which proved the rate $\mathcal{O}(n^{-1/3}\log^{2}n)$ for $\mathcal{H}^{\alpha}(R)$ with $\alpha>(d+4)/2$.
4 Binary classification
In binary classification, we observe a dataset $\mathcal{D}_{n}:=\{({\boldsymbol{X}}_{i},Y_{i}):i=1,\dots,n\}$ of $n$ i.i.d. copies of a random vector $({\boldsymbol{X}},Y)$, where we assume that the input vector ${\boldsymbol{X}}\in[0,1]^{d}$ and the label $Y\in\{-1,1\}$. The marginal distribution of ${\boldsymbol{X}}$ is denoted by $\mathbb{P}_{\boldsymbol{X}}$ and the conditional class probability function is denoted by
$\eta({\boldsymbol{x}}):=\mathbb{P}(Y=1|{\boldsymbol{X}}={\boldsymbol{x}}).$
For any real-valued function $f$ defined on $[0,1]^{d}$, we can define a classifier $\mathcal{C}_{f}({\boldsymbol{x}}):=\,{\rm sgn}\,(f({\boldsymbol{x}}))$. The classification error of $f$ is defined as
$\mathcal{E}(f)=\mathbb{E}_{{\boldsymbol{X}},Y}[\mathcal{C}_{f}({\boldsymbol{X}})\neq Y]=\mathbb{E}_{{\boldsymbol{X}},Y}[\boldsymbol{1}(Yf({\boldsymbol{X}})<0)],$
where $\boldsymbol{1}(\cdot)$ is 111 if ()($\cdot$)( ⋅ ) is true, and is 0 otherwise. A Bayes classifier $\mathcal{C}^{*}=\mathcal{C}_{f^{*}}$ is a classifier that minimizes the classification error $\mathcal{E}(f^{*})=\min_{f\in\mathcal{M}}\mathcal{E}(f)$, where $\mathcal{M}$ is the set of all measurable functions on $[0,1]^{d}$. Note that $\mathcal{C}^{*}=\,{\rm sgn}\,(2\eta-1)$ is a Bayes classifier and $\mathcal{E}(\mathcal{C}^{*})=\frac{1}{2}\mathbb{E}[1-|2\eta-1|]$. The goal of binary classification is to construct a classifier with small classification error by using the dataset $\mathcal{D}_{n}$.
Since we only have finite observed samples, one natural approach to estimate the Bayes classifier is the empirical risk minimization (with 010-10 - 1 loss)
(13) $\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{1}(Y_{i}f({\boldsymbol{X}}_{i})<0),$
where $\mathcal{F}_{n}$ is a prescribed function class. However, this procedure is often computational infeasible due to the NP-hardness of the minimization problem. In general, one replaces the 010-10 - 1 loss by surrogate losses. For a given surrogate loss function $\phi:\mathbb{R}\to[0,\infty)$, the $\phi$-risk is defined as
$\mathcal{L}_{\phi}(f)$ $:=\mathbb{E}_{{\boldsymbol{X}},Y}[\phi(Yf({\boldsymbol{X}}))]$
$=\mathbb{E}_{\boldsymbol{X}}[\eta({\boldsymbol{X}})\phi(f({\boldsymbol{X}}))+(1-\eta({\boldsymbol{X}}))\phi(-f({\boldsymbol{X}}))].$
Its minimizer is denoted by $f^{*}_{\phi}\in\operatorname*{argmin}_{f\in\mathcal{M}}\mathcal{L}_{\phi}(f)$. Note that $f^{*}_{\phi}$ can be explicitly computed by using the conditional class probability function $\eta$ for many convex loss functions $\phi$ [40, 35]. Instead of using (13), we can estimate the Bayes classifier by minimizing the empirical $\phi$-risk over a function class $\mathcal{F}_{n}$:
(14) $\widehat{f}_{\phi,n}\in\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}\phi(Y_{i}f({\boldsymbol{X}}_{i})).$
The goal of this section is to estimate the convergence rates of the excess classification risk and excess $\phi$-risk defined by
$\mathcal{R}(\widehat{f}_{\phi,n})$ $:=\mathcal{E}(\widehat{f}_{\phi,n})-\mathcal{E}(\mathcal{C}^{*}),$
$\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})$ $:=\mathcal{L}_{\phi}(\widehat{f}_{\phi,n})-\mathcal{L}_{\phi}(f^{*}_{\phi}),$
when $\mathcal{F}_{n}$ is parameterized by a CNN. The convergence rates certainly depend on properties of the conditional class probability function $\eta$. One of the well known assumptions on $\eta$ is the Tsybakov noise condition [21, 34]: there exist $q\in[0,\infty]$ and $c_{q}>0$ such that for any $t>0$,
(15) $\mathbb{P}_{\boldsymbol{X}}(|2\eta({\boldsymbol{X}})-1|\leq t)\leq c_{q}t^{q}.$
The constant $q$ is usually called the noise exponent. It is obvious that the Tsybakov noise condition always holds for $q=0$, whereas noise exponent $q=\infty$ means that $\eta$ is bounded away from the critical level 1/21/21 / 2. We will consider classifications with hinge loss and logistic loss under the Tsybakov noise condition.
4.1 Hinge loss
For the hinge loss $\phi(t)=\max\{1-t,0\}$, we have $f^{*}_{\phi}=\,{\rm sgn}\,(2\eta-1)=\mathcal{C}^{*}$ and $\mathcal{L}_{\phi}(f^{*}_{\phi})=\mathbb{E}[1-|2\eta-1|]$. It is well known that the following calibration inequality holds [40, 4]
(16) $\mathcal{R}(f)\leq\mathcal{R}_{\phi}(f).$
Hence, any convergence rate for the excess $\phi$-risk $\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})$ implies the same convergence rate for the excess classification risk $\mathcal{R}(\widehat{f}_{\phi,n})$. One can also check that [40, Section 3.3], if $|f|\leq 1$, then
(17) $\mathcal{R}_{\phi}(f)=\mathbb{E}[|f-f^{*}_{\phi}||2\eta-1|].$
To use this equality, it is natural to truncate the output of the estimator by using the truncation operator $\pi_{1}$ defined by (12).
In the following theorem, we provide convergence rates for the excess $\phi$-risk of the CNN classifier with hinge loss, under the assumption that the conditional class probability function $\eta$ is smooth and satisfies the Tsybakov noise condition.
Theorem 4.1.
Assume the noise condition (15) holds for some $q\in[0,\infty]$ and $\eta\in\mathcal{H}^{\alpha}(R)$ for some $0<\alpha<(d+3)/2$ and $R>0$. Let $\phi$ be the hinge loss and $\widehat{f}_{\phi,n}$ be the estimator defined by (14) with $\mathcal{F}_{n}=\{\pi_{1}f:f\in\mathcal{CNN}(s,J,L_{n},M_{n})\}$, where $s\in[2:d]$, $J\geq 6$ and
$L_{n}\asymp\left(\frac{n}{\log^{2}n}\right)^{\frac{d}{(q+2)\alpha+d}},\quad\left(\frac{n}{\log^{2}n}\right)^{\frac{3d+3}{2(q+2)\alpha+2d}}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n),$
then, for sufficiently large $n$,
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\right]\lesssim\left(\frac{\log^{2}n}{n}\right)^{\frac{(q+1)\alpha}{(q+2)\alpha+d}}.$
It was shown in [1] that the minimax lower bound for the excess classification risk is
(18) $\inf_{\widehat{f}_{n}}\sup_{\eta}\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}(\widehat{f}_{n})\right]\gtrsim n^{-\frac{(q+1)\alpha}{(q+2)\alpha+d}},$
where the supremum is taken over all $\eta\in\mathcal{H}^{\alpha}(R)$ that satisfies Tsybakov noise condition (15) and the infimum is taken over all estimators based on the training data $\mathcal{D}_{n}$. Hence, by the calibration inequality (16), the convergence rate in Theorem 4.1 is minimax optimal up to a logarithmic factor. Similar results have been established in [13] for fully connected neural networks with hinge loss. However, their results rely on the sparsity of neural networks and hence one need to optimize over different network architectures to obtain the optimal rate, which is hard to implement due to the unknown locations of the non-zero parameters. Our result show that CNNs, whose architecture is specifically defined, are able to achieve the optimal rate.
4.2 Logistic loss
For the logistic loss $\phi(t)=\log(1+e^{-t})$, we have $f^{*}_{\phi}=\log(\frac{\eta}{1-\eta})$ and $\mathcal{L}_{\phi}(f^{*}_{\phi})=\mathbb{E}[-\eta\log\eta-(1-\eta)\log(1-\eta)]$. Consequently, one can show that
$\mathcal{R}_{\phi}(f)=\mathbb{E}\left[\eta\log\left(\eta(1+e^{-f})\right)+(1-\eta)\log\left((1-\eta)(1+e^{f})\right)\right].$
Let us denote the KL-divergence by
$\mathcal{D}_{KL}(p_{1},p_{2}):=p_{1}\log\left(\frac{p_{1}}{p_{2}}\right)+(1-p_{1})\log\left(\frac{1-p_{1}}{1-p_{2}}\right),\quad p_{1},p_{2}\in[0,1],$
where $\mathcal{D}_{KL}(p_{1},p_{2})=\infty$ if $p_{2}=0$ and $p_{1}\neq 0$, or $p_{2}=1$ and $p_{1}\neq 1$. If we define the logistic function by
(19) $\psi(t):=\frac{1}{1+e^{-t}}\in[0,1],\quad t\in[-\infty,\infty],$
then a direct calculation shows that $\eta=\psi(f_{\phi}^{*})$ and
(20) $\mathcal{R}_{\phi}(f)=\mathbb{E}[\mathcal{D}_{KL}(\eta,\psi(f))].$
When the Tsybakov noise condition (15) holds, we have the following calibration inequality [32, Theorem 8.29]
(21) $\mathcal{R}(f)\leq 4c_{q}^{\frac{1}{q+2}}\mathcal{R}_{\phi}(f)^{\frac{q+1}{q+2}}.$
For the logistic loss, the convergence rate depends not only on the Tsybakov noise condition, but also upon the Small Value Bound (SVB) condition introduced by [5]. We say the distribution of $({\boldsymbol{X}},Y)$ satisfies the SVB condition, if there exists $\beta\geq 0$ and $C_{\beta}>0$ such that for any $t\in(0,1]$,
(22) $\mathbb{P}_{\boldsymbol{X}}(\eta({\boldsymbol{X}})\leq t)\leq C_{\beta}t^{\beta},\quad\mathbb{P}_{\boldsymbol{X}}(1-\eta({\boldsymbol{X}})\leq t)\leq C_{\beta}t^{\beta}.$
Note that this condition always holds for $\beta=0$ with $C_{\beta}=1$. The index $\beta$ is completely determined by the behavior of $\eta$ near 0 and 111. If $\eta$ is bounded away form 0 and 111, then the SVB condition holds for all $\beta>0$. In contrast, the Tsybakov noise condition provides a control on the behavior of $\eta$ near the decision boundary $\{{\boldsymbol{x}}:\eta({\boldsymbol{x}})=1/2\}$. This difference is due to the loss: the 010-10 - 1 loss only cares about the classification error, while the logistic loss measures how well the conditional class probability is estimated in the KL-divergence (20), which puts additional emphasis on small and large conditional class probabilities.
The following theorem gives convergence rates for CNNs under the SVB condition. As pointed out by [5], we do not get any gain in the convergence rate when the SVB index $\beta>1$. So, we assume $\beta\in[0,1]$ in the theorem.
Theorem 4.2.
Assume the SVB condition (22) holds for some $\beta\in[0,1]$ and $\eta\in\mathcal{H}^{\alpha}(R)$ for some $0<\alpha<(d+3)/2$ and $R>0$. Let $\phi$ be the logistic loss and $\widehat{f}_{\phi,n}$ be the estimator defined by (14) with $\mathcal{F}_{n}=\{\pi_{B_{n}}f:f\in\mathcal{CNN}(s,J,L_{n},M_{n})\}$, where $s\in[2:d]$, $J\geq 6$ and
$L_{n}\asymp\left(\frac{n}{\log n}\right)^{\frac{d}{(1+\beta)\alpha+d}},\quad\left(\frac{n}{\log n}\right)^{\frac{3d+3+2\alpha}{2(1+\beta)\alpha+2d}}\lesssim M_{n}\lesssim\,{\rm Poly}\,(n),\quad B_{n}\asymp\log n,$
then, for sufficiently large $n$,
$\mathbb{E}_{\mathcal{D}_{n}}\left[\mathcal{R}_{\phi}(\widehat{f}_{\phi,n})\right]\lesssim\left(\frac{\log n}{n}\right)^{\frac{(1+\beta)\alpha}{(1+\beta)\alpha+d}}\log n.$
The convergence rate in Theorem 4.2 is the same as [5, Theorem 3.3], which studied multi-class classification using fully-connected deep neural networks with cross entropy loss. If, in addition, the Tsybakov noise condition (15) holds, by combining Theorem 4.2 with the calibration inequality (21), we can get the following convergence rate for the classification risk:
$\mathbb{E}_{\mathcal{D}_{n}}[\mathcal{R}(\widehat{f}_{\phi,n})]\lesssim n^{-\frac{q+1}{q+2}\frac{(1+\beta)\alpha}{(1+\beta)\alpha+d}}\log^{2}n.$
Note that this rate is the same as the optimal rate (18) up to a logarithmic factor, when $q=0$ and $\beta=1$. For $\beta=0$, the obtained rate is not minimax optimal for the excess classification risk. However, as shown by [41, Corollary 2.1], the rate in Theorem 4.2 is indeed minimax optimal up to a logarithmic factor for the excess $\phi$-risk when $\beta=0$. So, even if the logistic classification can achieve the minimax optimal convergence rates for classification, it is in general not possible to derive it through the rates for excess $\phi$-risk.
There are other papers [14, 19, 26] studying the convergence rates of CNNs with logistic loss. The paper [14] imposed a max-pooling structure for the conditional class probability that is related to the structure of convolutional networks. So, their result is not comparable to ours. [19] used a similar setting as ours and derived the rate $n^{-\frac{\alpha}{2\alpha+2(\alpha\lor d)}}$ (ignoring logarithmic factors) for the excess $\phi$-risk under the assumption that $\eta$ is supported on a manifold of $d$ dimension. The article [26] also considered low-dimensional distributions and obtained the rate $n^{-\frac{\alpha}{2\alpha+d}}$ for the excess $\phi$-risk when $f^{*}_{\phi}\in\mathcal{H}^{\alpha}$, which is more restricted than the assumption $\eta\in\mathcal{H}^{\alpha}$. Our learning rate $n^{-\frac{\alpha}{\alpha+d}}$ in Theorem 4.2 (for $\beta=0$) is better than those in [19] and [26], but their results can be applied to low-dimensional distributions. It would be interesting to generalize our result to these distributions.
5 Conclusion
In this paper, we have studied approximation and learning capacities of convolutional neural networks with one-side zero-padding and multiple channels. We have derived new approximation bounds for CNNs with norm constraint on the weights. To study the generalization performance of learning algorithms induced by these networks, we also proved new bounds for their covering number. Based on these results, we established rates of convergence for CNNs in nonparametric regression and classification problems. Many of the obtained convergence rates are known to be minimax optimal.
There is a restriction on the smoothness of the target functions in our results. We think this restriction is due to the proof techniques of our approximation bound (Theorem 2.1), rather than the architecture of CNNs. It may be possible to use the ideas of network constructions from related works, such as [24, 42], to remove the restriction, which we leave as a future work.
Acknowledgments
This paper was submitted while Y. Yang was with City University of Hong Kong. The work described in this paper was partially supported by Discovery Project (DP240101919) of the Australian Research Council, InnoHK initiative, The Government of the HKSAR, Laboratory for AI-Powered Financial Technologies, the Research Grants Council of Hong Kong [Projects Nos. CityU 11315522, CityU 11303821] and National Natural Science Foundation of China [Project No. 12371103]. We thank the referees for their helpful comments and suggestions on the paper.